### v1

# Permutation-Invariant and Orientation-Aware Dataset Distillation for 3D Point Clouds

Abstract
We should collect large amount of data to train deep neural networks for various applications. Recently, the dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset while preserving essential task-relevant information. However, 3D point clouds distillation is almost unexplored due to the challenges of unordered structures of points. In this paper, we propose a novel distribution matching-based dataset distillation method for 3D point clouds that jointly optimizes the geometric structures of synthetic dataset as well as the orientations of synthetic models. To ensure the consistent feature alignment between different 3D point cloud models, we devise a permutation invariant distribution matching loss with the sorted feature vectors. We also employ learnable rotation angles to transform each syntheic model according to the optimal orientation best representing the original feature distribution. Extensive experimental results on widely used four benchmark datasets, including ModelNet10, ModelNet40, ShapeNet, and ScanObjectNN, demonstrate that the proposed method consistently outperforms the existing methods.
1 Introduction
With the increasing demand for large-scale datasets, dataset distillation [21] has gained significant attention across various applications of deep learning. The goal of dataset distillation is to reduce the size of an original dataset into a synthetic dataset that preserves the essential task-relevant information of the original dataset. Therefore, the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset. Existing dataset distillation methods [3, 6, 8, 11, 29, 28, 30, 17, 26, 5, 9] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches. The gradient matching and trajectory matching methods aim to ensure that the synthetic dataset produces similar optimization dynamics to the original dataset. To this end, the gradient matching method [29] minimizes the difference between the gradients computed on the synthetic and original datasets. However, these methods potentially overlook long-term dependencies in the training process. Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on a synthetic dataset follow similar optimization trajectories to those trained on the original dataset. However, these methods train the networks while optimizing the synthetic dataset significantly increasing the computational complexity. To alleviate the computational burden in dataset distillation, the distribution matching techniques [28, 30, 17, 26, 5] have been introduced. They randomly initialize the networks without training to extract the features, and compare the feature distributions between the original and synthetic datasets.
Figure 1: The concept of the proposed dataset distillation method for 3D point clouds. (a) Permutation-invariant feature mathcing. (b) Orientation optimization.
While the dataset distillation has been extensively studied for structured data, such as images [29, 3, 28] and texts [10, 12, 13], that for 3D point clouds is almost unexplored. Recently, PCC [27] applied an existing gradient matching method of DC [29] to tackle the data distillation problem of 3D point clouds, but it did not consider the inherent characteristics of 3D point clouds and still suffers from the high computational cost. Unlike structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space. Specifically, inconsistent indexing of points makes it challenging to develop efficient dataset distillation techniques. Furthermore, 3D models often exhibit different orientation from one another, that also degrades the distillation performance.
In this paper, we propose a joint optimization framework of Permutation Invariant Distribution Matching (PIDM) and orientation optimization for dataset disllation of 3D point clouds. A major challenge in this setting is achieving effective feature alignment despite the unordered nature of point clouds and their arbitrary orientations. As illustrated in Figure 1(a), the PIDM loss addresses the issue of unordered structure by sorting the elements within each feature vector according to their sizes before computing the distance between two feature distributions. To further enhance the feature alignment, as shown in Figure 1(b), We also employ learnable parameters of the rotation angles to estimate optimal orientation of 3D models, while optimizing the synthetic dataset. Experimental results demonstrate that the proposed method optimizes both the synthetic dataset and the orientation of 3D models simultaneously outperforming the existing methodologies.
The key contributions of this paper are summarized as follows:
  • To the best of our knowledge, we are the first to proposed a distribution matching-based dataset distillation method for 3D point clouds, that jointly optimizes the synthetic dataset and the orientation of 3D models simultaneously.
  • We devised PIDM loss to ensure permutation invariant feature alignment for 3D point clouds with inconsistent orders of points.
  • We validated proposed method through comprehensive experiments on four benchmark datasets of 3D point clouds classification, and showed the superiority of the proposed method over the existing dataset distillation techniques.
2 Related Works
Figure 2: The overall framework of the proposed 3D point clouds dataset distillation method.
2.1 3D Point Data Analysis
3D point clouds are generally unordered exhibiting irregular characteristics, that makes it difficult to apply the convolution operations in deep neural networks commonly used for images. Early approaches [25, 32] convert the point clouds into structured representation of multi-view images or voxel grids to enable the use of conventional deep learning architectures, such as convolutional neural networks (CNNs). VoxelNet [32] voxelizes the point clouds, but it causes increased memory consumption and quantization errors. To directly process point clouds without conversion to intermediate structured representations, PointNet [14] was introduced as a pioneering work that learns the features directly from unordered raw point clouds. However, it suffers from the limitation in capturing local geometric relationships, and hierarchical CNNs have been developed tailored to 3D point clouds. PointNet++[15] captures local geometric structures through iterative sampling and grouping strategies. On the other hand, PointConv [23] generalizes the convolution operations to irregular point clouds by adaptively weighting the neighboring points based on their spatial distribution. Meanwhile, DGCNN [22] constructs dynamic graphs from point clouds to adaptively capture the semantic information. Attempts have been also made to apply the transformer [20] to 3D point clouds processing, where the Point Transformer [31] utilizes the attention mechanism to capture the long-range dependencies.
2.2 Coreset Selection
Coreset selection is a technique designed to select representative samples from a given dataset, while maintaining the model’s performance even with the sampled data. The random selection method [16] randomly chooses a subset of data samples from the whole dataset. It is simple but suffers from the robustness due to the lack of informative criteria for sampling. The K-center method [18] selects data points iteratively that maximize the minimum distance to the set of already selected ones, taking into account the data distribution. The Herding method [2, 1] iteratively takes data points that minimize the discrepancy between the mean embeddings of the selected subset and the entire dataset in the feature space.
2.3 Dataset Distillation
Dataset distillation [21] methods can be largely categorized into the gradient matching [29], trajectory matching [3, 6, 8, 11], and distribution matching [28, 30, 17, 26, 5] approaches. The gradient matching, first introduced by DC [29], minimizes the difference between the gradients computed from the original and synthetic datasets, respectively, guiding the synthetic dataset to follow the training direction of the original dataset. The trajectory matching methods [3, 6, 8, 11], initially proposed by MTT [3], encourage the models trained on the synthetic dataset to follow the optimization trajectories similar to those trained on the original dataset rather than comparing individual gradients. ATT [11] automatically adjusts the length of the training trajectories between the synthetic and original datasets, enabling more effective and precise matching. The distribution matching, introduced by DM [28], focuses on minimizing the distance between the feature distributions of the original and synthetic datasets. DataDAM [17] enhances the distribution matching by aligning the feature maps using attention mechanism, achieving unbiased feature representation with low computational overhead. Furthermore, M3D [26] employs the Gaussian kernel function in the distribution matching loss, enabling the alignment across higher-order statistical characteristics of the feature distributions. Recently, a gradient matching-based point clouds distillation method [27] has been introduced, however it simply applied the existing image-based method without considering the unique characteristics of 3D point clouds.
3 Methodology
We propose a novel dataset distillation method for 3D point clouds that first addressing the key challenges of unordered structure and rotational variation of 3D point clouds. We perform PIDM that effectively aligns the features of unordered points across different 3D models. To handle the rotational variation, we also estimate optimal orientation while generating the synthetic dataset. Figure 2 shows the overall framework of the proposed method.
3.1 Problem Formulation
The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}$, $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, where the models trained on ${\boldsymbol{\mathcal{S}}}$ achieve close performance to that trained on ${\boldsymbol{\mathcal{T}}}$. Given a 3D point cloud sample ${{\bf p}}$ following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset ${\boldsymbol{\mathcal{S}}}^{\star}$ can be obtained via
${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}\limits_{{\boldsymbol{\mathcal{S}}}}\mathbb{E}_{{{\bf p}}}\left[||{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{T}}}}}({{\bf p}}),l)-{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{S}}}}}({{\bf p}}),l)||^{2}\right],$ (1)
where ${\mathcal{L}}$ denotes a task-specific loss function, such as the cross-entropy loss, and $\phi_{{\boldsymbol{\mathcal{T}}}}$ and $\phi_{{\boldsymbol{\mathcal{S}}}}$ represent the models trained on ${\boldsymbol{\mathcal{T}}}$ and ${\boldsymbol{\mathcal{S}}}$, respectively.
3.2 Permutation Invariant Distribution Matching
We basically perform the distribution matching by minimizing the discrepancy between the feature distributions derived from the original and synthetic datasets, respectively, such that
${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}_{{\boldsymbol{\mathcal{S}}}}D(\phi({\boldsymbol{\mathcal{T}}}),\phi({\boldsymbol{\mathcal{S}}})),$ (2)
where $D$ is a distance function, and $\phi$ denotes a randomly initialized untrained network. Note that we simply use a randomly initialized network without training and effectively alleviate the computational overhead in dataset distillation. The Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used to measure $D$, defined as
$\mathcal{L}_{\text{MMD}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}}})-2{K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}}),$ (3)
where ${K}(\cdot,\cdot)$ is the kernel function, where we used the Gaussian kernel in this work, given by
${K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\phi({{\bf t}})-\phi({{\bf s}})||^{2}}{2\sigma}\right).$ (4)
Note that the spatial relationships among the pixels are well structured and consistent across different images, but the points in 3D space are usually indexed in different orders across multiple 3D models. Therefore, the conventional distribution matching methods cannot be directly applied to 3D point clouds, since they assume aligned features between two compared models. To overcome this issue, we propose PIDM that performs the feature sorting in advance when comparing the feature distributions between the original and synthetic datasets. Specifically, given a 3D point cloud model ${{\bf p}}$, we extract point-wise features by using $\phi$ as
$\phi({{\bf p}})=\left[{\bf f}_{1},{\bf f}_{2},\dots,{\bf f}_{C}\right],\quad{\bf f}_{i}\in\mathbb{R}^{N},$ (5)
where $N$ denotes the number of points in ${{\bf p}}$, and $C$ is the number of feature channels. We first rearrange the elements of ${\bf f}_{i}$ in descending order of their sizes to have the sorted feature vector $\tilde{{\bf f}}_{i}$. Then we have the sorted features as
$\tilde{\phi}({{\bf p}})=\left[\tilde{{\bf f}}_{1},\tilde{{\bf f}}_{2},\dots,\tilde{{\bf f}}_{C}\right],\quad\tilde{{\bf f}}_{i}\in\mathbb{R}^{N}.$ (6)
Note that ${{\bf t}}$ in the original dataset $\boldsymbol{\mathcal{T}}$ and ${{\bf s}}$ in the synthetic dataset $\boldsymbol{\mathcal{S}}$ exhibit different orders of points in general, and therefore the features in ${\phi}({{\bf t}})$ and ${\phi}({{\bf s}})$ are inconsistenly aligned with each other disrupting the desired behavior of comparison in (4). On the contrary, $\tilde{\phi}({{\bf t}})$ and $\tilde{\phi}({{\bf s}})$ yield well aligned features invariant to the permutation variation, and therefore, we compute the refined Gaussian kernel with the sorted features, given by
$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
Accordingly, we devise the PIDM loss $\mathcal{L}_{\text{PIDM}}$ from $\mathcal{L}_{\text{MMD}}$, given by
$\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}}})-2\tilde{K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}}).$ (8)
By using $\mathcal{L}_{\text{PIDM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we improve the performance of datasets distillation for 3D point clouds regardless of the orders of points.
Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
PPC 1 3 10 1 3 10 1 3 10 1 3 10
Ratio (%) 0.25 0.75 2.5 0.4 1.2 4.0 0.15 0.45 1.5 0.15 0.45 1.5
Whole 91.41 87.84 82.49 63.84
Random 35.52 75.17 85.31 34.57 59.96 74.06 33.96 54.84 63.13 13.85 20.42 34.75
Herding 40.07 78.02 86.86 54.38 68.45 78.84 49.49 59.79 66.85 15.67 27.66 38.67
K-center 40.07 77.59 83.22 54.38 63.03 65.26 49.49 51.39 47.82 15.67 19.81 24.00
DM 31.86 77.56 86.10 32.29 61.97 75.23 27.84 56.06 64.27 14.89 23.25 35.58
DC 43.89 74.97 86.11 52.00 66.56 75.57 49.78 60.08 64.58 18.63 24.30 35.27
Ours 44.70 84.96 87.79 55.80 72.08 80.07 50.20 63.74 68.35 17.29 31.84 43.91
Table 1: Performance comparison between three coreset selection methods [16, 2, 18], DC [29], DM [28], and the proposed method. DC, DM, and proposed method were initialized with random selection [16] for fair comparison. The best and the second best scores are highlighted in bold and underlined, respectively.
3.3 Orientation Optimization
3D models usually exhibit different orientations with each other. Therefore, while generating optimal 3D models in the synthetic dataset in terms of the geometric shapes, we also estimate their optimal orientations best representing the various orientations of 3D models in the original dataset. In practice, we employ three rotation angles of $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, around the $x$, $y$, and $z$-axes, respectively, as learnable parameters which are used to adaptively adjust the orientation of each 3D model in the synthetic dataset ${\boldsymbol{\mathcal{S}}}$ during the dataset distillation. Therefore, instead of $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (8), we minimize $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset where its 3D models are rotated toward the estimated optimal orientation.
Note that we optimize the synthetic 3D models as well as their orientations simultaneously, during the dataset distillation process. Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset. We then have $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the 3D models in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to 0. Then, for each class, a mini-batch is sampled from the original dataset $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per each class. The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with its batch size set to the number of point clouds per class (PPC). The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and the orientation parameters of its models by minimizing $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$. This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while optimizing the orientation.
4 Experimental Results
4.1 Experimental Setup
Datasets. The proposed method was evaluated on the ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively. The ScanObjectNN consists of real-world scanned objects from 15 classes, and hence a more challenging dataset due to the presence of noise, occlusions, and other artifacts commonly found in scanned 3D models.
Evaluation Metrics. Following previous methods [29, 28], we measure the classification accuracy trained on the distilled synthetic dataset. Since the training dynamics of the network can vary with each run, we ensure the fairness by training the network 10 times and report the mean accuracy across these runs. Similar to Images Per Class (IPC) in the previous image-based distillation, we evaluate the performance across different Point clouds Per Class (PPC) values. Experiments were conducted under the PPC settings of 1, 3, and 10 to analyze the impact of varying data compression levels.
Implementation Details. Each point cloud model contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15]. The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iteration by training the PointNet on the synthetic dataset and testing it on the original test dataset. Additionally, SGD was used as the optimizer for $\boldsymbol{\mathcal{S}}$, with a learning rate of 10. For optimizing the rotation angles, SGD was also used, with a learning rate of 0.5 for $\theta_{x}$ and $\theta_{z}$. Since the most datasets are aligned with the y-axis, a higher learning rate of 5 was applied for optimizing $\theta_{y}$. PointNet was optimized using the SGD optimizer with a learning rate of 0.01, and the training was conducted for 500 iterations. Also, note that we only use the distribution of the features before the max pooling to compute the PIDM loss to retain fine-grained structural information. To further emphasize the most dominant features after sorting, the loss computed from the largest magnitude values is also used.
4.2 Performance Comparison
We compare the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18]. We additionally compare DC [29] and DM [28], representative distillation methods for 2D images, to provide a broader perspective on dataset distillation. We also compare PCC [27], the only existing 3D point clouds dataset distillation method.
Table 1 compares the quantitative performance of the proposed method with that of the existing methodologies. DM and DC were initialized with random selection [16]. The synthetic datasets were optimized using PointNet [14], and classification performance was evaluated on PointNet to demonstrate their validity. Among the coreset selection methods, Herding [2, 1] achieves the best overall performance. When PPC is set to 1, Herding and K-Center [18] yield identical results since K-Center selects the first data point using the same algorithm to Herding. As PPC increases, Herding consistently outperforms other coreset selection methods, reflecting its effectiveness. However, since coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions.
DM [28] matches the distribution of the global features extracted by PointNet. However, PointNet applies the global max pooling before classification, retaining only the most dominant signal per channel while discarding substantial information. This loss of information weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset. DC [29], on the other hand, tends to fit the specific network used for distillation due to its structure designed for learning the network. However, PointNet [14], which is used for distillation, is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients. As a result, it consistently underperforms compared to the proposed method. Furthermore, as the PPC increases, the optimization process involves a larger number of point clouds, making the gradient alignment increasingly unstable and leading to reduced performance.
The proposed method outperforms both the coreset selection and existing distillation techniques. When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement. As the PPC increases, the method more effectively matches the feature distributions significantly improving the performance.
Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
Method DM DC Ours DM DC Ours DM DC Ours DM DC Ours
PointNet++ 34.97 51.49 82.44 14.14 46.88 68.45 20.26 34.30 53.20 15.45 11.57 19.70
DGCNN 56.60 62.58 79.01 51.65 52.30 66.86 13.77 37.98 52.81 17.72 12.49 19.02
PointConv 33.41 39.96 56.57 20.00 37.90 51.32 17.35 33.99 47.41 15.16 14.20 16.82
PT 48.52 61.81 77.62 26.70 47.54 61.95 34.15 42.62 55.48 17.06 14.89 21.87
Table 2: Comparison of cross-architecture generalization performance of DC [29], DM [28], and the proposed method, evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
Init Method MN10 MN40 SN SONN
Noise DC 18.77 6.05 5.24 12.73
DM 26.53 12.04 12.09 13.55
Ours 70.69 65.25 58.44 29.03
Random DC 68.32 64.71 58.15 26.07
DM 65.17 56.50 49.39 24.57
Ours 72.48 69.31 60.76 31.01
K-center DC 69.57 61.85 42.88 21.65
DM 63.04 49.93 44.82 21.59
Ours 72.72 68.20 60.70 28.25
Herding DC 71.50 66.79 59.23 26.98
DM 64.76 59.16 52.61 27.58
Ours 72.93 69.67 62.25 31.08
Table 3: Performance comparison across different initialization strategies including uniform noise, random selection [16], Herding [2], and K-center [18]. We averaged the accuracies of thre cases of PPC, 1, 3, and 10. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
4.2.1 Cross-Architecture Generalization
We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31]. Unlike PointNet, the architectures used for generalization performance comparison follow a hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from PointNet. This difference makes it essential for a distilled dataset to retain rich and transferable features that generalize well across architectures.
Since the use of symmetric functions like global max pooling causes significant information loss, DM fails to preserve the original feature distribution and instead learns only the features relevant to PointNet. This limitation prevents the synthetic dataset from capturing diverse structural details, making it highly specific to PointNet. As a result, when evaluated on hierarchical models, DM performs poorly, as it lacks the necessary information for effective generalization. DC, while performing better than DM, also suffers from key limitations. Its gradient matching approach forces the synthetic dataset to align with whole gradients of PointNet parameters, which makes it more effective within the same architecture but limits its adaptability to other models. Also, the bi-level optimization process further aggravates this issue by overfitting the synthetic dataset to the specific training conditions of PointNet. Consequently, when applied to other architectures, DC struggles to generalize effectively.
The proposed method addresses these limitations by direct feature distribution matching with randomly initialized network instead of bi-level optimization. By comparing feature maps with richer information, the synthetic dataset captures features relevant to PointNet while following the original data distribution more closely. This broader feature matching strategy improves performance across different architectures. Unlike DC, which suffers from overfitting, and DM, which loses critical information, the proposed approach adapts more effectively, showing strong generalization across various backbone networks.
Method MN10 MN40 SN SONN
DC 1.52 5.43 7.57 2.19
DM 0.04 0.11 0.15 0.05
Ours 0.08 0.26 0.35 0.11
Table 4: Comparison of training time (in hours) of DC, DM, and the proposed method. Experiments were conducted at PPC 3. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
Figure 3: Comparison of synthetic datasets distilled from the ModelNet40 [24] dataset. Point clouds were colorized according to the y-coordinates. From top to bottom, the orginal dataset (first), initialized point cloud models (second), and the distilled synthetic datasets by using DC [29] (third), DM [28] (fourth), and the proposed method (fifth).
MN10 MN40 SN SONN
Random 75.17 59.96 54.84 20.42
w/o sorting 75.01 59.96 53.91 20.20
w/ sorting 84.96 72.08 63.74 31.57
Table 5: Ablation study of the proposed sorting method. Experiments were performed at PPC 3. “Random” denotes the results of the random selection [16]. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
4.2.2 Initialization Strategies
Table 3 presents the classification performance on PointNet for DC [29], PCC [27] and DM [28] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18], compared to the proposed method. Note that the results of PCC are obtained from our implementation, where PCC refers to performing DC on a dataset initialized using Herding.
When initialized with noise, DC fails to effectively optimize the synthetic dataset, as the lack of geometric information disrupts the gradient computation. As a result, the original dataset fails to guide the training process along a meaningful path, leading to poor convergence. In contrast, structured initializations like random selection, Herding, or K-Center improve performance, with Herding performing best by selecting the most representative samples. Unlike DC, DM does not include network training process, making it less sensitive to noise initialization and allowing for better performance in such cases. However, when initialized with random selection, K-Center, or Herding, DM consistently underperforms DC due to its reliance on feature representations that discard significant information.
Unlike DC or DM, whose dataset distillation algorithms are unstable, the proposed method achieves comparable performance even when initialized from uniform noise. Furthermore, the proposed method significantly outperforms other approaches under the heuristic initialization strategy, such as random selection. This demonstrates that the proposed method is robust to initialization and effectively resolves the instability issues inherent in previous dataset distillation approaches.
${\boldsymbol{\mathcal{S}}}$ ${\boldsymbol{\mathcal{\theta}}}$ Aligned Mixed Rotated
_ _ 65.01 59.96 20.42
_ 66.22 62.17 24.89
_ 73.96 71.51 29.72
74.35 72.08 31.84
Table 6: Ablation study of optimizing the synthetic dataset and orientation. Experiments were performed at PPC 3. The baseline (first row) shows the performance without optimization which is identical to that of the random selection.
4.2.3 Training Time Comparison
Table 4 compares the training time between the proposed and previous methods. DM [28] achieves the fastest training time as it uses only a small subset of the overall features. However, this limited feature usage causes many useful features to be ignored, leading to performance degradation. In contrast, DC [29] incurs significantly higher computational cost as it aligns gradients across all parameters. Notably, for the largest dataset, ShapeNet [4], DC requires over 7 hours of training, whereas DM completes training in just 9 minutes (0.15 hours), making DC more than 50 times slower. This latency is even more pronounced in point cloud datasets, as networks designed for point cloud processing, such as PointNet [14], have higher computational complexity than those used in image dataset distillation. Since our algorithm is based on distribution matching approach, it requires only a marginal increase in training time compared to the DM algorithm, yet achieves a significant improvement in performance.
Layer 1 Layer 2 Layer 3 PointNet
_ _ 61.47
_ _ 62.22
_ _ 62.72
Table 7: Ablation study of the impact of different PointNet layers. The performance was averaged across the four models of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. Experiments were performed at PPC 3.
4.3 Visualization
Figure 3 presents a visualization of the distilled synthetic dataset generated by previous methods and the proposed method. DC [29], which relies on gradient matching, struggles to effectively match the gradients of the large PointNet [14] network. As a result, it fails to deviate significantly from the initialized point cloud model, leading to minimal movement in certain points while introducing noise. DM [28], on the other hand, performs feature distribution matching after the global max pooling operation in PointNet, which loses critical information. As a result, similar to DC, DM preserves the original shape while only shifting specific points, failing to capture meaningful structural changes. In contrast, the proposed method successfully preserves the semantic structure of each class while selectively learning essential features. Notably, in the airplane class, the wing structure appears distinct compared to other methods. Additionally, in certain categories, such as guitar, both the shape and orientation are changed, demonstrating that the proposed method functions as intended. These results highlight the effectiveness of our approach in producing a cleaner, noise-free synthetic dataset with superior visual quality compared to previous methods.
4.4 Ablation Study
4.4.1 Effect of Permutation Invariant Distribution Matching Loss
Table 5 presents an ablation study on the impact of the proposed sorting mechanism on classification performance using PointNet [14]. Without the sorting mechanism (w/o $\mathcal{F}$), the unordered nature of point clouds restricts effective feature matching, resulting in performance that is nearly identical to that of random selection [16], indicating that the model has not been properly optimized. Otherwise, when the sorting mechanism is applied (w/ $\mathcal{F}$), accuracy improves across all datasets, as the sorted feature provide a consistent and structured representation that facilitates more effective feature matching during optimization. This result confirms that addressing the inherent permutation-invariance of point clouds through sorting enables the network to align corresponding features more effectively, resulting in a substantial improvement in classification performance.
4.4.2 Effect of Orientation Optimization
Table 6 presents an ablation study to evaluate the impact of optimizing the synthetic dataset and rotation angles independently and jointly, using the proposed PIDM loss. We evaluated on dataset groups categorized into aligned, mixed, and rotated to analyze the effectiveness of the proposed orientation optimization. The aligned group consists of datasets where objects maintain a consistent orientation, including ModelNet10 and ShapeNet [4]. The mixed group includes datasets where only certain classes exhibit rotation variations, represented by ModelNet40 [24]. The rotated group comprises datasets in which objects exhibit arbitrary rotations across all classes, such as ScanObjectNN [19].
In the second row, we present the results where only the rotation angle is optimized without optimizing the synthetic dataset. Introducing the proposed orientation optimization method, optimized using the PIDM loss, improves performance across all groups compared to the non-optimized setting. The improvement becomes more noticeable as the class with rotations becomes more dominant in the dataset, indicating the importance of orientation optimization in determining the optimal direction for data with arbitrary rotations. The addition of synthetic dataset optimization further boosts performance, with the best results achieved when both optimizations are applied jointly. Specifically, the accuracy on the rotated dataset increases significantly, showing that jointly optimizing the synthetic dataset and rotation angles with the PIDM loss helps the model better adapt to rotation variations.
4.4.3 Effect of Feature Extraction Layers in PointNet
Table 7 presents an ablation study examining how classification performance on PointNet [14] varies when using feature maps from different layers. PointNet consists of three layers before the pooling operation, which produce feature maps with 64, 128, and 1024 channels, respectively. The results show that deeper feature extraction layers generally lead to better performance. This performance improvement can be attributed to the increased number of channels, which allows the feature map to capture more detailed and diverse information. A higher-dimensional representation provides greater capacity to maintain essential features from the original dataset, enabling the network to better approximate the underlying distribution.
5 Conclusion
We proposed a permutation-invariant and orientation-aware dataset distillation framework for 3D point clouds. To alleviate the inconsistency of point ordering between compared 3D models, we devise a Permutation Invariant Distribution Matching (PIDM) loss where the sorted features are compared. We also employed learnable parameters of rotation angles to estimate optimal orientation of synthetic models. The geometric structures and orientation of the synthetic models are jointly optimized during the dataset distillation. Experimental results evaluated on four widely used benchmark datasets of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] demonstrate that the proposed method outperforms existing distillation methods while maintaining strong cross-architecture generalization. Furthermore, we also validated the effectiveness of each component of the proposed method through extensive ablation studies.

### v2

@@
-Permutation-Invariant and Orientation-Aware Dataset Distillation for 3D Point Clouds
+Dataset Distillation of 3D Point Clouds via Distribution Matching
 Abstract
-We should collect large amount of data to train deep neural networks for various applications.
-Recently, the dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset while preserving essential task-relevant information.
-However, 3D point clouds distillation is almost unexplored due to the challenges of unordered structures of points.
-In this paper, we propose a novel distribution matching-based dataset distillation method for 3D point clouds that jointly optimizes the geometric structures of synthetic dataset as well as the orientations of synthetic models.
-To ensure the consistent feature alignment between different 3D point cloud models, we devise a permutation invariant distribution matching loss with the sorted feature vectors.
-We also employ learnable rotation angles to transform each syntheic model according to the optimal orientation best representing the original feature distribution.
-Extensive experimental results on widely used four benchmark datasets, including ModelNet10, ModelNet40, ShapeNet, and ScanObjectNN, demonstrate that the proposed method consistently outperforms the existing methods.
+Large-scale datasets are usually required to train deep neural networks, but it increases the computational complexity hindering the practical applications.
+Recently, dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset to alleviate the computational burden of training while preserving essential task-relevant information.
+However, the dataset distillation for 3D point clouds remains largely unexplored, as the point clouds exhibit fundamentally different characteristics from that of images, making the dataset distillation more challenging.
+In this paper, we propose a distribution matching-based distillation framework for 3D point clouds that jointly optimizes the geometric structures as well as the orientations of the synthetic 3D objects.
+To address the semantic misalignment caused by unordered indexing of points, we introduce a Semantically Aligned Distribution Matching loss computed on the sorted features in each channel.
+Moreover, to address the rotation variation, we jointly learn the optimal rotation angles while updating the synthetic dataset to better align with the original feature distribution.
+Extensive experiments on widely used benchmark datasets demonstrate that the proposed method consistently outperforms existing dataset distillation methods, achieving superior accuracy and strong cross-architecture generalization.
 1 Introduction
-With the increasing demand for large-scale datasets, dataset distillation [21] has gained significant attention across various applications of deep learning.
-The goal of dataset distillation is to reduce the size of an original dataset into a synthetic dataset that preserves the essential task-relevant information of the original dataset.
-Therefore, the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset.
+With the increasing demand for large-scale training datasets, the cost of model retraining from scratch has become a major challenge leading to the research of dataset distillation [21].
+The objective of dataset distillation is to produce a significantly smaller synthetic dataset from a large original dataset, that preserves the essential task-relevant information contained in the original dataset.
+Hence the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset.
 Existing dataset distillation methods [3, 6, 8, 11, 29, 28, 30, 17, 26, 5, 9] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches.
@@
 To this end, the gradient matching method [29] minimizes the difference between the gradients computed on the synthetic and original datasets.
-However, these methods potentially overlook long-term dependencies in the training process.
-Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on a synthetic dataset follow similar optimization trajectories to those trained on the original dataset.
-However, these methods train the networks while optimizing the synthetic dataset significantly increasing the computational complexity.
+However, it potentially overlooks long-term dependencies in the training process.
+Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on the synthetic dataset follow similar optimization trajectories to those trained on the original dataset.
+Note that these methods train the networks while optimizing the synthetic dataset, and significantly increase the computational complexity.
 To alleviate the computational burden in dataset distillation, the distribution matching techniques [28, 30, 17, 26, 5] have been introduced.
 They randomly initialize the networks without training to extract the features, and compare the feature distributions between the original and synthetic datasets.
-Figure 1: The concept of the proposed dataset distillation method for 3D point clouds. (a) Permutation-invariant feature mathcing. (b) Orientation optimization.
-While the dataset distillation has been extensively studied for structured data, such as images [29, 3, 28] and texts [10, 12, 13], that for 3D point clouds is almost unexplored.
-Recently, PCC [27] applied an existing gradient matching method of DC [29] to tackle the data distillation problem of 3D point clouds, but it did not consider the inherent characteristics of 3D point clouds and still suffers from the high computational cost.
-Unlike structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space.
-Specifically, inconsistent indexing of points makes it challenging to develop efficient dataset distillation techniques.
-Furthermore, 3D models often exhibit different orientation from one another, that also degrades the distillation performance.
-In this paper, we propose a joint optimization framework of Permutation Invariant Distribution Matching (PIDM) and orientation optimization for dataset disllation of 3D point clouds.
+While dataset distillation has been extensively studied for structured data such as images [29, 3, 28] and texts [10, 12, 13], its application to 3D point clouds remains almost unexplored.
+PCC [27] simply applied an existing distillation method for images [29] to the 3D point clouds without considering inherent characteristics of 3D point clouds, however it still suffers from high computational complexity.
+Unlike the structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space.
+Therefore, semantically similar regions across different 3D models are often associated with inconsistent orders (indices of points), that makes direct feature comparison incorrect worsening the performance of distribution matching based dataset distillation.
+We refer to this issue as semantic misalignment.
+Moreover, the datasets of 3D point clouds also suffers from rotational variation.
+3D point clouds are often captured or synthesized under arbitrary poses due to the absence of canonical orientations.
+Such rotational difference causes the objects of the same class yield different features from one another, significantly increasing the intra-class variability.
+As a result, it becomes difficult to construct a representative synthetic dataset that faithfully follows the feature distribution of the original dataset.
+In this paper, we propose an optimization framework combining Semantically Aligned Distribution Matching (SADM) and orientation optimization for dataset distillation of 3D point clouds.
 A major challenge in this setting is achieving effective feature alignment despite the unordered nature of point clouds and their arbitrary orientations.
-As illustrated in Figure 1(a), the PIDM loss addresses the issue of unordered structure by sorting the elements within each feature vector according to their sizes before computing the distance between two feature distributions.
-To further enhance the feature alignment, as shown in Figure 1(b), We also employ learnable parameters of the rotation angles to estimate optimal orientation of 3D models, while optimizing the synthetic dataset.
-Experimental results demonstrate that the proposed method optimizes both the synthetic dataset and the orientation of 3D models simultaneously outperforming the existing methodologies.
+The SADM loss addresses the issue of misaligned semantic structures by sorting the point-wise feature values within each channel before computing the distance between two feature distributions.
+In parallel, we employ learnable rotation parameters to address the orientation variation by estimating the optimal poses of 3D models while generating the synthetic dataset.
+By simultaneously enforcing semantic consistency and orientation alignment, the proposed method achieves superior performance of dataset distillation compared with the existing methods, as demonstrated by extensive experiments on standard 3D point cloud classification benchmarks.
 The key contributions of this paper are summarized as follows:
-  • To the best of our knowledge, we are the first to proposed a distribution matching-based dataset distillation method for 3D point clouds, that jointly optimizes the synthetic dataset and the orientation of 3D models simultaneously.
-  • We devised PIDM loss to ensure permutation invariant feature alignment for 3D point clouds with inconsistent orders of points.
-  • We validated proposed method through comprehensive experiments on four benchmark datasets of 3D point clouds classification, and showed the superiority of the proposed method over the existing dataset distillation techniques.
+  • To the best of our knowledge, we are the first to propose a distribution matching-based dataset distillation method of 3D point clouds, that jointly optimizes the shapes and orientations of synthetic dataset.
+  • We devised the SADM loss on the sorted features in each channel to preserve the semantic alignment between compared 3D objects.
+  • We validated the proposed method through extensive experiments on the four benchmark datasets widely used for 3D point clouds classification, and showed the superiority of the proposed method over the existing dataset distillation techniques.
 2 Related Works
-Figure 2: The overall framework of the proposed 3D point clouds dataset distillation method.
 2.1 3D Point Data Analysis
@@
 3 Methodology
-We propose a novel dataset distillation method for 3D point clouds that first addressing the key challenges of unordered structure and rotational variation of 3D point clouds.
-We perform PIDM that effectively aligns the features of unordered points across different 3D models.
-To handle the rotational variation, we also estimate optimal orientation while generating the synthetic dataset.
+Figure 1: The overall framework of the proposed dataset distillation method for 3D point clouds.
+We propose a novel dataset distillation method for 3D point clouds that first addresses the key challenges of semantic misalignment and rotational variation.
+We perform SADM that effectively aligns the features of semantically consistent structures across different 3D models.
+To handle the rotational variation, we also estimate optimal orientations while generating the synthetic dataset.
 Figure 1 shows the overall framework of the proposed method.
-3.1 Problem Formulation
-The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}$, $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, where the models trained on ${\boldsymbol{\mathcal{S}}}$ achieve close performance to that trained on ${\boldsymbol{\mathcal{T}}}$.
+3.1 Preliminaries
+Problem Definition.
+The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}_{i=1}^{|\mathcal{T}|}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}_{i=1}^{|\mathcal{S}|}$, where $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, such that a model trained on ${\boldsymbol{\mathcal{S}}}$ achieves close performance to that trained on ${\boldsymbol{\mathcal{T}}}$.
 Given a 3D point cloud sample ${{\bf p}}$ following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset ${\boldsymbol{\mathcal{S}}}^{\star}$ can be obtained via
 ${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}\limits_{{\boldsymbol{\mathcal{S}}}}\mathbb{E}_{{{\bf p}}}\left[||{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{T}}}}}({{\bf p}}),l)-{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{S}}}}}({{\bf p}}),l)||^{2}\right],$ (1)
-where ${\mathcal{L}}$ denotes a task-specific loss function, such as the cross-entropy loss, and $\phi_{{\boldsymbol{\mathcal{T}}}}$ and $\phi_{{\boldsymbol{\mathcal{S}}}}$ represent the models trained on ${\boldsymbol{\mathcal{T}}}$ and ${\boldsymbol{\mathcal{S}}}$, respectively.
-3.2 Permutation Invariant Distribution Matching
-We basically perform the distribution matching by minimizing the discrepancy between the feature distributions derived from the original and synthetic datasets, respectively, such that
+where ${\mathcal{L}}$ denotes a task-specific loss function, such as the cross-entropy loss, and $\phi_{{\boldsymbol{\mathcal{T}}}}$ and $\phi_{{\boldsymbol{\mathcal{S}}}}$ represent the models to estimate the class label which are trained on ${\boldsymbol{\mathcal{T}}}$ and ${\boldsymbol{\mathcal{S}}}$, respectively.
+Distribution Matching. Distribution matching (DM) strategy focuses on aligning the feature distributions derived from the original and synthetic datasets, respectively, via
 ${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}_{{\boldsymbol{\mathcal{S}}}}D(\phi({\boldsymbol{\mathcal{T}}}),\phi({\boldsymbol{\mathcal{S}}})),$ (2)
-where $D$ is a distance function, and $\phi$ denotes a randomly initialized untrained network.
-Note that we simply use a randomly initialized network without training and effectively alleviate the computational overhead in dataset distillation.
-The Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used to measure $D$, defined as
+where $\phi$ is the feature extractor and $D$ denotes a distance function.
+We employ a randomly initialized, untrained network, which has been demonstrated to sufficiently capture the structural information for distribution alignment [28].
+Also, the Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used as $D$, given by
 $\mathcal{L}_{\text{MMD}}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}}},{\boldsymbol{\mathcal{S}}})-2{K}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{S}}}),$ (3)
-where ${K}(\cdot,\cdot)$ is the kernel function, where we used the Gaussian kernel in this work, given by
-${K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\phi({{\bf t}})-\phi({{\bf s}})||^{2}}{2\sigma}\right).$ (4)
-Note that the spatial relationships among the pixels are well structured and consistent across different images, but the points in 3D space are usually indexed in different orders across multiple 3D models.
-Therefore, the conventional distribution matching methods cannot be directly applied to 3D point clouds, since they assume aligned features between two compared models.
-To overcome this issue, we propose PIDM that performs the feature sorting in advance when comparing the feature distributions between the original and synthetic datasets.
-Specifically, given a 3D point cloud model ${{\bf p}}$, we extract point-wise features by using $\phi$ as
+where ${K}(\cdot,\cdot)$ is a kernel function. We used the Gaussian kernel in this work as
+${K}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{\bf t}}\in\boldsymbol{\mathcal{T}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\exp\left(-\frac{\|\phi({{\bf t}})-\phi({{\bf s}})\|^{2}}{2\sigma}\right).$ (4)
+3.2 Semantically Aligned Distribution Matching
+Figure 2: Visualization of the points corresponding to the largest (top), 200th largest (middle), and 500th largest (bottom) features in each channel.
+Whereas the pixels of image exhibit well structured spatial relationships with one another and consistently indexed across different images, the points in 3D point clouds are unordered with different indices across different models.
+Therefore, the conventional distribution matching methods cannot be directly applied to 3D point clouds, since the features of semantically similar structures are not aligned between two compared models.
+Motivated by the importance of preserving semantic correspondence, we investigate the relationship between the feature values and semantic significance in 3D point clouds.
+We first extracted point-wise features using a randomly initialized network $\phi$, and sorted the feature values for each of 1024 channels according to their size.
+Figure 2 visualizes the points corresponding to the largest (top), 200th largest (middle), and 500th largest (bottom) features, respectively.
+We observe that, even without model training, the points associated with the largest features consistently capture semantically meaningful regions, such as edges and corners, across different classes.
+In contrast, the points yielding low-ranked features are distributed around less informative regions.
+This suggests that the features reflect the relative importance of points in characterizing the structures of 3D objects, and the points with similar orders of sorted features tend to capture semantically related regions across different 3D models.
+In order to preserve the semantic correspondence across different 3D models, we propose a semantically aligned distribution matching (SADM) loss.
+Specifically, given a 3D point cloud object ${{\bf p}}$, we extract the features using $\phi$ via
 $\phi({{\bf p}})=\left[{\bf f}_{1},{\bf f}_{2},\dots,{\bf f}_{C}\right],\quad{\bf f}_{i}\in\mathbb{R}^{N},$ (5)
-where $N$ denotes the number of points in ${{\bf p}}$, and $C$ is the number of feature channels.
-We first rearrange the elements of ${\bf f}_{i}$ in descending order of their sizes to have the sorted feature vector $\tilde{{\bf f}}_{i}$.
-Then we have the sorted features as
+where $N$ denotes the number of points in ${{\bf p}}$, $C$ is the number of feature channels, and ${\bf f}_{i}=\left\{f_{i,1},f_{i,2},...,f_{i,N}\right\}$ represents the feature vector of the $i$-th channel where $f_{i,j}$ is the feature value of the $j$-th point.
+We then perform channel-wise sorting to the extracted features of $f_{i,j}$’s in the descending order of their sizes, and obtain the sorted feature vector $\tilde{{\bf f}}_{i}=\left\{\tilde{f}_{i,1},\tilde{f}_{i,2},...,\tilde{f}_{i,N}\right\}$ such that $\tilde{f}_{i,j}\geq\tilde{f}_{i,j+1}$.
+Then we have the set of the sorted feature vectors as
 $\tilde{\phi}({{\bf p}})=\left[\tilde{{\bf f}}_{1},\tilde{{\bf f}}_{2},\dots,\tilde{{\bf f}}_{C}\right],\quad\tilde{{\bf f}}_{i}\in\mathbb{R}^{N}.$ (6)
-Note that ${{\bf t}}$ in the original dataset $\boldsymbol{\mathcal{T}}$ and ${{\bf s}}$ in the synthetic dataset $\boldsymbol{\mathcal{S}}$ exhibit different orders of points in general, and therefore the features in ${\phi}({{\bf t}})$ and ${\phi}({{\bf s}})$ are inconsistenly aligned with each other disrupting the desired behavior of comparison in (4).
-On the contrary, $\tilde{\phi}({{\bf t}})$ and $\tilde{\phi}({{\bf s}})$ yield well aligned features invariant to the permutation variation, and therefore, we compute the refined Gaussian kernel with the sorted features, given by
-$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
-Accordingly, we devise the PIDM loss $\mathcal{L}_{\text{PIDM}}$ from $\mathcal{L}_{\text{MMD}}$, given by
-$\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}}})-2\tilde{K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}}).$ (8)
-By using $\mathcal{L}_{\text{PIDM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we improve the performance of datasets distillation for 3D point clouds regardless of the orders of points.
+Note that ${{\bf t}}$ in the original dataset $\boldsymbol{\mathcal{T}}$ and ${{\bf s}}$ in the synthetic dataset $\boldsymbol{\mathcal{S}}$ exhibit different orders of points in general, and therefore the features in ${\phi}({{\bf t}})$ and ${\phi}({{\bf s}})$ are inconsistently aligned with each other worsening the desired behavior of comparison in (3).
+On the contrary, the sorted features $\tilde{\phi}({{\bf t}})$ and $\tilde{\phi}({{\bf s}})$ exhibit consistent ordering of features in each channel reflecting their relative semantic importance, and thus facilitates reliable feature comparison across different 3D objects.
+To measure the discrepancy between the sorted features, we redefine the Gaussian kernel as
+$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\exp\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
+Then we devise the loss $\mathcal{L}_{\alpha}$ from $\mathcal{L}_{\text{MMD}}$ (3), given by
+$\mathcal{L}_{\alpha}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})=\tilde{K}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}})+\tilde{K}(\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}})-2\tilde{K}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}).$ (8)
+We additionally employ $\mathcal{L}_{\beta}$ to boost the role of the most significant feature in each channel, defined as
+$\mathcal{L}_{\beta}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})=\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}})+\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}})-2\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (9)
+where $\tilde{K}_{\text{top}}(\cdot,\cdot)$ denotes the Gaussian kernel computed with only the most largest feature in each channel.
+The SADM loss is then formulated as a weighted sum of the two losses, given by
+$\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})=\lambda_{1}\mathcal{L}_{\alpha}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})+\lambda_{2}\mathcal{L}_{\beta}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (10)
+where $\lambda_{1},\lambda_{2}$ are the weighting parameters.
+By using $\mathcal{L}_{\text{SADM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we facilitate reliable matching of feature distributions considering sematic structures of 3D point clouds, thereby improving the performance dataset distillation.
+3.3 Estimation of Optimal Rotations
+3D objects usually exhibit different orientations from each other.
+Therefore, while generating optimal 3D objects in the synthetic dataset in terms of their geometric shapes, we also estimate their optimal rotations best representing various orientations of 3D objects in the original dataset.
+In practice, we introduce three rotation angles, $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, corresponding to rotations around the $x$-, $y$-, and $z$-axes, respectively.
+These angles are treated as learnable parameters, allowing the orientation of each synthetic 3D object in ${\boldsymbol{\mathcal{S}}}$ to be adaptively adjusted during dataset distillation.
+Therefore, instead of minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (10), we minimize $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset whose 3D objects are rotated toward the estimated optimal orientations.
+Note that we optimize the synthetic 3D objects as well as their rotation parameters simultaneously, during the dataset distillation process.
+Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset.
+We then form $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the objects in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to zero.
+For each class, a mini-batch is sampled from $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per class.
+The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with the batch size determined by the number of point cloud objects per class (PPC).
+The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and their rotation parameters by minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$.
+This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while aligning their orientations more effectively.
+The benefit of this joint optimization is justified in the following proposition.
+Proposition 1.
+Jointly optimizing the synthetic dataset $\boldsymbol{\mathcal{S}}$ and the rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ guarantees a lower or equal loss to that of optimizing $\boldsymbol{\mathcal{S}}$ alone.
+$\min_{\boldsymbol{\mathcal{S}},\boldsymbol{\theta}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}}))\leq\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (11)
+where $\mathcal{R}_{\boldsymbol{\theta}}$ denotes the rotation operator according to $\boldsymbol{\theta}$.
+Proof Sketch.
+The proof is provided in Appendix A.
+Allowing optimization over both $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ enlarges the feasible set compared to optimizing only $\boldsymbol{\mathcal{S}}$.
+The feasible set for the joint optimization includes the feasible set of the original optimization as a subset, since optimizing only $\boldsymbol{\mathcal{S}}$ is equivalent to the special case of the joint optimization with the fixed $(\theta_{x},\theta_{y},\theta_{z})=(0,0,0)$.
+Consequently, the minimum loss achieved by jointly optimizing over $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ must be less than or equal to the minimum loss achieved by optimizing only $\boldsymbol{\mathcal{S}}$.
+$\square$
+Table 1: Comparison of quantitative performance.
+DC, DM, and the proposed method were initialized with random selection [16] for fair comparison.
+The best and the second best scores are highlighted in bold and underlined, respectively.
 Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
@@
 Whole 91.41 87.84 82.49 63.84
-Random 35.52 75.17 85.31 34.57 59.96 74.06 33.96 54.84 63.13 13.85 20.42 34.75
-Herding 40.07 78.02 86.86 54.38 68.45 78.84 49.49 59.79 66.85 15.67 27.66 38.67
-K-center 40.07 77.59 83.22 54.38 63.03 65.26 49.49 51.39 47.82 15.67 19.81 24.00
-DM 31.86 77.56 86.10 32.29 61.97 75.23 27.84 56.06 64.27 14.89 23.25 35.58
-DC 43.89 74.97 86.11 52.00 66.56 75.57 49.78 60.08 64.58 18.63 24.30 35.27
-Ours 44.70 84.96 87.79 55.80 72.08 80.07 50.20 63.74 68.35 17.29 31.84 43.91
-Table 1: Performance comparison between three coreset selection methods [16, 2, 18], DC [29], DM [28], and the proposed method.
-DC, DM, and proposed method were initialized with random selection [16] for fair comparison.
-The best and the second best scores are highlighted in bold and underlined, respectively.
-3.3 Orientation Optimization
-3D models usually exhibit different orientations with each other.
-Therefore, while generating optimal 3D models in the synthetic dataset in terms of the geometric shapes, we also estimate their optimal orientations best representing the various orientations of 3D models in the original dataset.
-In practice, we employ three rotation angles of $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, around the $x$, $y$, and $z$-axes, respectively, as learnable parameters which are used to adaptively adjust the orientation of each 3D model in the synthetic dataset ${\boldsymbol{\mathcal{S}}}$ during the dataset distillation.
-Therefore, instead of $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (8), we minimize $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset where its 3D models are rotated toward the estimated optimal orientation.
-Note that we optimize the synthetic 3D models as well as their orientations simultaneously, during the dataset distillation process.
-Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset.
-We then have $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the 3D models in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to 0.
-Then, for each class, a mini-batch is sampled from the original dataset $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per each class.
-The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with its batch size set to the number of point clouds per class (PPC).
-The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and the orientation parameters of its models by minimizing $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$.
-This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while optimizing the orientation.
+Random 35.5±$\pm 4.7$ 75.2±$\pm 1.7$ 85.3±$\pm 1.1$ 34.6±$\pm 1.8$ 60.0±$\pm 1.32$ 74.1±$\pm 0.4$ 34.0±$\pm 3.0$ 54.8±$\pm 1.5$ 63.1±$\pm 1.0$ 13.9±$\pm 1.4$ 20.4±$\pm 1.3$ 34.8±$\pm 1.1$
+Herding 40.1 ±$\pm 5.2$ 78.0 ±$\pm 1.3$ 86.9±$\pm 0.6$ 54.4 ±$\pm 2.0$ 68.5 ±$\pm 1.0$ 78.8 ±$\pm 0.4$ 49.5 ±$\pm 2.3$ 59.8 ±$\pm 0.9$ 66.9 ±$\pm 0.5$ 15.7 ±$\pm 1.7$ 27.7 ±$\pm 1.4$ 38.7 ±$\pm 1.7$
+K-center 40.1 ±$\pm 5.2$ 77.6±$\pm 1.9$ 83.2±$\pm 1.4$ 54.4±$\pm 2.0$ 63.0 ±$\pm 2.7$ 65.3 ±$\pm 1.1$ 49.5 ±$\pm 2.3$ 51.4 ±$\pm 1.7$ 47.8 ±$\pm 0.7$ 15.7 ±$\pm 1.7$ 19.8 ±$\pm 0.8$ 24.0 ±$\pm 1.0$
+DM 31.9±$\pm 4.3$ 77.6±$\pm 1.6$ 86.1±$\pm 0.8$ 32.3±$\pm 4.8$ 62.0±$\pm 1.9$ 75.2±$\pm 0.5$ 27.8±$\pm 4.0$ 56.1±$\pm 1.2$ 64.3±$\pm 0.7$ 14.9±$\pm 2.0$ 23.3±$\pm 1.5$ 35.6±$\pm 1.4$
+DC 43.9±$\pm 5.5$ 75.0±$\pm 2.4$ 86.1±$\pm 1.1$ 52.0±$\pm 1.7$ 66.6±$\pm 1.3$ 75.6±$\pm 0.5$ 49.8±$\pm 1.3$ 60.1±$\pm 1.4$ 64.6±$\pm 0.7$ 18.6±$\pm\textbf{1.6}$ 24.3±$\pm 2.5$ 35.3±$\pm 1.2$
+Ours 44.7 ±$\pm\textbf{6.1}$ 84.4 ±$\pm\textbf{1.2}$ 87.8 ±$\pm\textbf{1.0}$ 55.8 ±$\pm\textbf{1.5}$ 72.1 ±$\pm\textbf{0.8}$ 80.1 ±$\pm\textbf{0.4}$ 50.2 ±$\pm\textbf{1.4}$ 63.7 ±$\pm\textbf{0.7}$ 68.4 ±$\pm\textbf{0.5}$ 17.4 ±$\pm 1.5$ 31.6 ±$\pm\textbf{1.0}$ 43.9 ±$\pm\textbf{1.9}$
+Table 2: Comparison of cross-architecture generalization performance evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
+Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
+Method DM DC Ours DM DC Ours DM DC Ours DM DC Ours
+PointNet++ 35.0±$\pm 8.3$ 51.5±$\pm 7.8$ 82.9±$\pm\textbf{1.4}$ 14.1±$\pm 2.7$ 46.9±$\pm 6.0$ 68.5±$\pm\textbf{0.5}$ 20.3±$\pm 3.7$ $\pm 3.4$ 53.2±$\pm\textbf{1.5}$ 15.5±$\pm 3.6$ 11.6±$\pm 2.2$ 22.0±$\pm\textbf{3.9}$
+DGCNN 56.6±$\pm 5.3$ 62.6±$\pm 2.2$ 79.0±$\pm\textbf{2.2}$ 34.8±$\pm 3.5$ 52.3±$\pm 2.7$ 66.9±$\pm\textbf{1.1}$ 13.8±$\pm 2.2$ 38.0±$\pm 2.7$ 52.8±$\pm\textbf{1.6}$ 17.7±$\pm 2.4$ 12.5±$\pm 1.3$ 19.0±$\pm\textbf{1.5}$
+PointConv 33.4±$\pm 6.2$ 40.0±$\pm 10.5$ 56.6±$\pm\textbf{4.9}$ 20.0±$\pm 5.7$ 37.9±$\pm 3.9$ 51.3±$\pm\textbf{6.2}$ 17.4±$\pm 3.0$ 34.0±$\pm 4.8$ 47.4±$\pm\textbf{2.3}$ 15.2±$\pm 1.7$ 14.2±$\pm 1.8$ 16.8±$\pm\textbf{1.5}$
+PT 48.5±$\pm 6.6$ $\pm 1.6$ 77.6±$\pm\textbf{1.6}$ 26.7±$\pm 5.0$ 47.5±$\pm 2.7$ 61.6±$\pm\textbf{1.1}$ 34.2±$\pm 5.3$ 42.6±$\pm 2.5$ 55.5±$\pm\textbf{0.5}$ 17.1±$\pm 0.8$ 14.9±$\pm 0.5$ 21.9±$\pm\textbf{2.7}$
 4 Experimental Results
 4.1 Experimental Setup
-Datasets.
 The proposed method was evaluated on the ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets.
 The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively.
-The ScanObjectNN consists of real-world scanned objects from 15 classes, and hence a more challenging dataset due to the presence of noise, occlusions, and other artifacts commonly found in scanned 3D models.
-Evaluation Metrics.
+The ScanObjectNN consists of real-world scanned objects from 15 classes.
 Following previous methods [29, 28], we measure the classification accuracy trained on the distilled synthetic dataset.
-Since the training dynamics of the network can vary with each run, we ensure the fairness by training the network 10 times and report the mean accuracy across these runs.
-Similar to Images Per Class (IPC) in the previous image-based distillation, we evaluate the performance across different Point clouds Per Class (PPC) values.
-Experiments were conducted under the PPC settings of 1, 3, and 10 to analyze the impact of varying data compression levels.
-Implementation Details.
-Each point cloud model contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15].
+We ensure the fairness by training the network 10 times and report the mean and standard deviation of accuracy.
+We evaluate the performance across different PPC values of 1, 3, and 10, respectively.
+Each point cloud object contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15].
 The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iteration by training the PointNet on the synthetic dataset and testing it on the original test dataset.
-Additionally, SGD was used as the optimizer for $\boldsymbol{\mathcal{S}}$, with a learning rate of 10.
-For optimizing the rotation angles, SGD was also used, with a learning rate of 0.5 for $\theta_{x}$ and $\theta_{z}$.
-Since the most datasets are aligned with the y-axis, a higher learning rate of 5 was applied for optimizing $\theta_{y}$.
-PointNet was optimized using the SGD optimizer with a learning rate of 0.01, and the training was conducted for 500 iterations.
-Also, note that we only use the distribution of the features before the max pooling to compute the PIDM loss to retain fine-grained structural information.
-To further emphasize the most dominant features after sorting, the loss computed from the largest magnitude values is also used.
 4.2 Performance Comparison
-We compare the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18].
-We additionally compare DC [29] and DM [28], representative distillation methods for 2D images, to provide a broader perspective on dataset distillation.
-We also compare PCC [27], the only existing 3D point clouds dataset distillation method.
-Table 1 compares the quantitative performance of the proposed method with that of the existing methodologies.
-DM and DC were initialized with random selection [16].
+4.2.1 Dataset Distillation
+We compare the performance of the proposed method with that of 1) the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18], 2) two existing image dataset distillation methods of DC [29] and DM [28], and 3) the only existing 3D point clouds dataset distillation method of PCC [27].
+Table 1 compares the quantitative performance of the proposed method with that of the existing methods.
+DM and DC were initialized with random selection.
 The synthetic datasets were optimized using PointNet [14], and classification performance was evaluated on PointNet to demonstrate their validity.
-Among the coreset selection methods, Herding [2, 1] achieves the best overall performance.
-When PPC is set to 1, Herding and K-Center [18] yield identical results since K-Center selects the first data point using the same algorithm to Herding.
-As PPC increases, Herding consistently outperforms other coreset selection methods, reflecting its effectiveness.
-However, since coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions.
-DM [28] matches the distribution of the global features extracted by PointNet.
-However, PointNet applies the global max pooling before classification, retaining only the most dominant signal per channel while discarding substantial information.
-This loss of information weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset.
-DC [29], on the other hand, tends to fit the specific network used for distillation due to its structure designed for learning the network.
-However, PointNet [14], which is used for distillation, is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients.
+Among the coreset selection methods, Herding achieves the best overall performance.
+When PPC is set to 1, Herding and K-Center yield identical results since K-Center selects the first data point using the same algorithm to Herding.
+As PPC increases, Herding consistently outperforms other coreset selection methods, showing its effectiveness.
+However, since the coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions.
+The features used in DM are extracted from the layer before the classifier, and they discard substantial information through the global max pooling process of PointNet.
+This information loss weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset.
+DC, on the other hand, aims to match the gradients of the network.
+However, PointNet is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients.
 As a result, it consistently underperforms compared to the proposed method.
-Furthermore, as the PPC increases, the optimization process involves a larger number of point clouds, making the gradient alignment increasingly unstable and leading to reduced performance.
 The proposed method outperforms both the coreset selection and existing distillation techniques.
 When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement.
-As the PPC increases, the method more effectively matches the feature distributions significantly improving the performance.
-Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
-Method DM DC Ours DM DC Ours DM DC Ours DM DC Ours
-PointNet++ 34.97 51.49 82.44 14.14 46.88 68.45 20.26 34.30 53.20 15.45 11.57 19.70
-DGCNN 56.60 62.58 79.01 51.65 52.30 66.86 13.77 37.98 52.81 17.72 12.49 19.02
-PointConv 33.41 39.96 56.57 20.00 37.90 51.32 17.35 33.99 47.41 15.16 14.20 16.82
-PT 48.52 61.81 77.62 26.70 47.54 61.95 34.15 42.62 55.48 17.06 14.89 21.87
-Table 2: Comparison of cross-architecture generalization performance of DC [29], DM [28], and the proposed method, evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
+As the PPC increases, the proposed method more effectively matches the feature distributions significantly improving the performance.
+Table 3: Performance comparison across different initialization strategies.
+The accuracies of three PPC values of 1, 3, and 10 are averaged.
+MN10: ModelNet10.
+MN40: ModelNet40.
+SN: ShapeNet.
+SONN: ScanObjectNN.
 Init Method MN10 MN40 SN SONN
@@
 Ours 72.93 69.67 62.25 31.08
-Table 3: Performance comparison across different initialization strategies including uniform noise, random selection [16], Herding [2], and K-center [18].
-We averaged the accuracies of thre cases of PPC, 1, 3, and 10.
-MN10: ModelNet10.
-MN40: ModelNet40.
-SN: ShapeNet.
-SONN: ScanObjectNN.
-4.2.1 Cross-Architecture Generalization
-We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31].
-Unlike PointNet, the architectures used for generalization performance comparison follow a hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from PointNet.
-This difference makes it essential for a distilled dataset to retain rich and transferable features that generalize well across architectures.
-Since the use of symmetric functions like global max pooling causes significant information loss, DM fails to preserve the original feature distribution and instead learns only the features relevant to PointNet.
-This limitation prevents the synthetic dataset from capturing diverse structural details, making it highly specific to PointNet.
-As a result, when evaluated on hierarchical models, DM performs poorly, as it lacks the necessary information for effective generalization.
-DC, while performing better than DM, also suffers from key limitations.
-Its gradient matching approach forces the synthetic dataset to align with whole gradients of PointNet parameters, which makes it more effective within the same architecture but limits its adaptability to other models.
-Also, the bi-level optimization process further aggravates this issue by overfitting the synthetic dataset to the specific training conditions of PointNet.
-Consequently, when applied to other architectures, DC struggles to generalize effectively.
-The proposed method addresses these limitations by direct feature distribution matching with randomly initialized network instead of bi-level optimization.
-By comparing feature maps with richer information, the synthetic dataset captures features relevant to PointNet while following the original data distribution more closely.
-This broader feature matching strategy improves performance across different architectures.
-Unlike DC, which suffers from overfitting, and DM, which loses critical information, the proposed approach adapts more effectively, showing strong generalization across various backbone networks.
+† DC [29] with Herding initialization corresponds to PCC [27]
+Table 4: Comparison of training time (in hours) at the PPC value of 3.
 Method MN10 MN40 SN SONN
@@
 Ours 0.08 0.26 0.35 0.11
-Table 4: Comparison of training time (in hours) of DC, DM, and the proposed method. Experiments were conducted at PPC 3. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
+Table 5: Ablation study of the proposed SADM loss.
+Experiments were performed at the PPC value of 3.
+“Random” denotes the results of the random selection [16], and SA is the proposed semantic alignment.
+MN10 MN40 SN SONN
+Random 75.17 59.96 54.84 20.42
+w/o SA 75.01 59.96 53.91 20.20
+w/ SA 84.42 72.08 63.74 31.84
+4.2.2 Cross-Architecture Generalization
+We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31], respectively.
+Unlike PointNet, the architectures used for generalization performance comparison follow hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from that of PointNet.
+This difference makes it essential for the distilled dataset to retain rich and transferable features that generalize well toward different architectures.
+Due to significant information loss caused by the pooling operation, DM primarily captures the features specific to the PointNet, failing to preserve the original feature distribution.
+Consequently, the synthetic dataset lacks diverse structural details, resulting in poor generalization performance, particularly on hierarchical models.
+DC, while performing better than DM, also suffers from key limitations.
+It matches the gradients obtained from a trained PointNet, the synthetic dataset becomes overfitted to that specific architecture and fails to generalize toward other networks.
+Consequently, the generalization ability of DC degrades when applied to other architectures.
+In contrary, the proposed method addresses these limitations by directly matching the feature distributions extracted from randomly initialized network.
+By comparing feature maps containing richer and less architecture-biased information, the synthetic dataset captures the features relevant to the PointNet while following the original data distribution more closely.
+Wherease DC suffers from the overfitting and DM loses critical information, the proposed method yields higher adaptability showing strong generalization performance across various backbone networks.
 Figure 3: Comparison of synthetic datasets distilled from the ModelNet40 [24] dataset.
@@
 From top to bottom, the orginal dataset (first), initialized point cloud models (second), and the distilled synthetic datasets by using DC [29] (third), DM [28] (fourth), and the proposed method (fifth).
-MN10 MN40 SN SONN
-Random 75.17 59.96 54.84 20.42
-w/o sorting 75.01 59.96 53.91 20.20
-w/ sorting 84.96 72.08 63.74 31.57
-Table 5: Ablation study of the proposed sorting method.
-Experiments were performed at PPC 3.
-“Random” denotes the results of the random selection [16].
-MN10: ModelNet10.
-MN40: ModelNet40.
-SN: ShapeNet.
-SONN: ScanObjectNN.
 4.2.3 Initialization Strategies
-Table 3 presents the classification performance on PointNet for DC [29], PCC [27] and DM [28] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18], compared to the proposed method.
-Note that the results of PCC are obtained from our implementation, where PCC refers to performing DC on a dataset initialized using Herding.
-When initialized with noise, DC fails to effectively optimize the synthetic dataset, as the lack of geometric information disrupts the gradient computation.
-As a result, the original dataset fails to guide the training process along a meaningful path, leading to poor convergence.
-In contrast, structured initializations like random selection, Herding, or K-Center improve performance, with Herding performing best by selecting the most representative samples.
-Unlike DC, DM does not include network training process, making it less sensitive to noise initialization and allowing for better performance in such cases.
-However, when initialized with random selection, K-Center, or Herding, DM consistently underperforms DC due to its reliance on feature representations that discard significant information.
-Unlike DC or DM, whose dataset distillation algorithms are unstable, the proposed method achieves comparable performance even when initialized from uniform noise.
+Table 5 presents the classification performance on PointNet for DC [29], DM [28], and PCC [27] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18].
+Note that the results of PCC are obtained by our implementation, where the PCC is equivalent to performing DC with Herding initialization.
+When initialized with noise, DC fails to effectively optimize the synthetic dataset because the network has far more parameters than typical networks used for image dataset distillation, making it difficult to align the gradients.
+Therefore, the original dataset cannot properly guide the training process, leading to poor convergence.
+In contrast, the structured initializations such as random selection, Herding, or K-Center improve the performance, where the Herding performs the best by selecting the most representative samples.
+Unlike DC, DM directly matches the feature representations rather than relying on the gradients and hence less sensitive to the initialization strategies.
+However, due to the information loss, DM consistently underperforms DC even with the structured initializations.
+In contrast, the proposed method achieves favorable performance, even when initialized from uniform noise.
 Furthermore, the proposed method significantly outperforms other approaches under the heuristic initialization strategy, such as random selection.
-This demonstrates that the proposed method is robust to initialization and effectively resolves the instability issues inherent in previous dataset distillation approaches.
+This demonstrates that the proposed method is robust with diverse initialization schemes and effectively resolves the instability issues inherent in the existing dataset distillation methods.
+4.2.4 Training Time
+Table 5 compares the training times between the proposed method and the existing methods.
+DM [28] achieves the fastest training time as it uses only a small subset of the overall features, however it often ignores many useful features leading to performance degradation.
+In contrast, DC [29] incurs a significantly higher computational cost as it aligns the gradients across all the parameters.
+Notably, for the largest dataset of ShapeNet [4], DC requires over 7 hours for training 50 times slower that DM that completes the training in just 9 minutes (0.15 hours).
+Since the proposed method is based on the distribution matching approach, its training time is marginally increased compared to DM, yet achieves significant improvement in dataset distillation performance.
+4.2.5 Qualitative Comparison of Synthetic Datasets
+Figure 3 compares the resulting synthetic datasets distilled by using the proposed method and the existing methods, respectively.
+The first row shows the original datasets and the second row shows initialized point cloud objects.
+DC [29] fails to deviate significantly from the initialized objects and causes noise.
+Similarly, DM [28] also maintains the original shape of the initialized objects while shifting certain points only, that hinders to capture meaningful structural changes.
+In contrast, as shown in the bottom row, the proposed method successfully preserves the overall semantic structures of 3D objects in each class, while selectively learning essential features.
+For example, in the airplane class, we observe significant changes in the edges and corners, that are critical for class discrimination, especially around the wings.
+Also, in the guitar class, the proposed method jointly optimizes the shape and orientation of the synthesized 3D objects.
+4.3 Ablation Study
+4.3.1 Effect of SADM Loss
+Table 5 presents the ablation study on the impact of the proposed SADM loss on the classification performance using PointNet [14].
+Without preserving the semantic alignment, the unordered nature of point clouds restricts effective feature matching, resulting in nearly identical performance to that of the random selection [16].
+That means the model has not been properly optimized.
+Otherwise, when the semantic alignment is applied, accuracy improves across all datasets, as the aligned featured provide consistent and structured representation that facilitates more effective feature matching during the optimization.
+This result confirms that preserving the inherent semantic alignment via feature sorting facilitates the network to exploit well-aligned corresponding features and significantly improves the classification performance.
+4.3.2 Effect of Optimal Rotation Estimation
+Table 6: Ablation study of the proposed optimal rotation estimation. Experiments were performed at the PPC value of 3.
 ${\boldsymbol{\mathcal{S}}}$ ${\boldsymbol{\mathcal{\theta}}}$ Aligned Mixed Rotated
@@
 74.35 72.08 31.84
-Table 6: Ablation study of optimizing the synthetic dataset and orientation.
-Experiments were performed at PPC 3.
+Table 6 presents the ablation study to evaluate the impact of joint optimization of synthetic dataset and rotation angles.
+We categorized the test datsets into the aligned, mixed, and rotated groups to analyze the effectiveness of the proposed rotation estimation.
+The aligned group consists of the datasets where the objects maintain consistent orientations, including ModelNet10 and ShapeNet [4].
+The mixed group includes ModelNet40 [24] where only certain classes exhibit rotation variations.
+The rotated group is ScanObjectNN [19] where the objects exhibit arbitrary rotations across all classes.
 The baseline (first row) shows the performance without optimization which is identical to that of the random selection.
-4.2.3 Training Time Comparison
-Table 4 compares the training time between the proposed and previous methods.
-DM [28] achieves the fastest training time as it uses only a small subset of the overall features.
-However, this limited feature usage causes many useful features to be ignored, leading to performance degradation.
-In contrast, DC [29] incurs significantly higher computational cost as it aligns gradients across all parameters.
-Notably, for the largest dataset, ShapeNet [4], DC requires over 7 hours of training, whereas DM completes training in just 9 minutes (0.15 hours), making DC more than 50 times slower.
-This latency is even more pronounced in point cloud datasets, as networks designed for point cloud processing, such as PointNet [14], have higher computational complexity than those used in image dataset distillation.
-Since our algorithm is based on distribution matching approach, it requires only a marginal increase in training time compared to the DM algorithm, yet achieves a significant improvement in performance.
-Layer 1 Layer 2 Layer 3 PointNet
-_ _ 61.47
-_ _ 62.22
-_ _ 62.72
-Table 7: Ablation study of the impact of different PointNet layers.
-The performance was averaged across the four models of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets.
-Experiments were performed at PPC 3.
-4.3 Visualization
-Figure 3 presents a visualization of the distilled synthetic dataset generated by previous methods and the proposed method.
-DC [29], which relies on gradient matching, struggles to effectively match the gradients of the large PointNet [14] network.
-As a result, it fails to deviate significantly from the initialized point cloud model, leading to minimal movement in certain points while introducing noise.
-DM [28], on the other hand, performs feature distribution matching after the global max pooling operation in PointNet, which loses critical information.
-As a result, similar to DC, DM preserves the original shape while only shifting specific points, failing to capture meaningful structural changes.
-In contrast, the proposed method successfully preserves the semantic structure of each class while selectively learning essential features.
-Notably, in the airplane class, the wing structure appears distinct compared to other methods.
-Additionally, in certain categories, such as guitar, both the shape and orientation are changed, demonstrating that the proposed method functions as intended.
-These results highlight the effectiveness of our approach in producing a cleaner, noise-free synthetic dataset with superior visual quality compared to previous methods.
-4.4 Ablation Study
-4.4.1 Effect of Permutation Invariant Distribution Matching Loss
-Table 5 presents an ablation study on the impact of the proposed sorting mechanism on classification performance using PointNet [14].
-Without the sorting mechanism (w/o $\mathcal{F}$), the unordered nature of point clouds restricts effective feature matching, resulting in performance that is nearly identical to that of random selection [16], indicating that the model has not been properly optimized.
-Otherwise, when the sorting mechanism is applied (w/ $\mathcal{F}$), accuracy improves across all datasets, as the sorted feature provide a consistent and structured representation that facilitates more effective feature matching during optimization.
-This result confirms that addressing the inherent permutation-invariance of point clouds through sorting enables the network to align corresponding features more effectively, resulting in a substantial improvement in classification performance.
-4.4.2 Effect of Orientation Optimization
-Table 6 presents an ablation study to evaluate the impact of optimizing the synthetic dataset and rotation angles independently and jointly, using the proposed PIDM loss.
-We evaluated on dataset groups categorized into aligned, mixed, and rotated to analyze the effectiveness of the proposed orientation optimization.
-The aligned group consists of datasets where objects maintain a consistent orientation, including ModelNet10 and ShapeNet [4].
-The mixed group includes datasets where only certain classes exhibit rotation variations, represented by ModelNet40 [24].
-The rotated group comprises datasets in which objects exhibit arbitrary rotations across all classes, such as ScanObjectNN [19].
-In the second row, we present the results where only the rotation angle is optimized without optimizing the synthetic dataset.
-Introducing the proposed orientation optimization method, optimized using the PIDM loss, improves performance across all groups compared to the non-optimized setting.
-The improvement becomes more noticeable as the class with rotations becomes more dominant in the dataset, indicating the importance of orientation optimization in determining the optimal direction for data with arbitrary rotations.
-The addition of synthetic dataset optimization further boosts performance, with the best results achieved when both optimizations are applied jointly.
-Specifically, the accuracy on the rotated dataset increases significantly, showing that jointly optimizing the synthetic dataset and rotation angles with the PIDM loss helps the model better adapt to rotation variations.
-4.4.3 Effect of Feature Extraction Layers in PointNet
-Table 7 presents an ablation study examining how classification performance on PointNet [14] varies when using feature maps from different layers.
-PointNet consists of three layers before the pooling operation, which produce feature maps with 64, 128, and 1024 channels, respectively.
-The results show that deeper feature extraction layers generally lead to better performance.
-This performance improvement can be attributed to the increased number of channels, which allows the feature map to capture more detailed and diverse information.
-A higher-dimensional representation provides greater capacity to maintain essential features from the original dataset, enabling the network to better approximate the underlying distribution.
+In the second row, we present the results where only the orientation of the synthetic data is optimized using the proposed SADM loss, without updating the synthetic dataset.
+This leads to performance improvements across all the groups compared to random selection.
+The improvement becomes more significant as the classes with rotation variation become more dominant.
+Optimizing the synthetic dataset further boosts the performance, with the best results achieved when both the synthetic data and rotation angles are optimized simultaneously.
+In particular, the accuracy on the rotated dataset improves significantly, demonstrating that the proposed joint optimization enables the model to be resilient to rotation variations.
+4.4 Limitation
+While the proposed SADM loss enhances the semantic consistency by sorting the features, it does not completely align semantically meaningful regions across different point cloud objects.
+Estimating optimal rotations also improves the complexity, particularly when the datasets are already well-aligned to the canonical axes.
 5 Conclusion
-We proposed a permutation-invariant and orientation-aware dataset distillation framework for 3D point clouds.
-To alleviate the inconsistency of point ordering between compared 3D models, we devise a Permutation Invariant Distribution Matching (PIDM) loss where the sorted features are compared.
-We also employed learnable parameters of rotation angles to estimate optimal orientation of synthetic models.
-The geometric structures and orientation of the synthetic models are jointly optimized during the dataset distillation.
-Experimental results evaluated on four widely used benchmark datasets of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] demonstrate that the proposed method outperforms existing distillation methods while maintaining strong cross-architecture generalization.
+We proposed a semantically-aligned and orientation-aware dataset distillation framework for 3D point clouds.
+To alleviate the inconsistency of point ordering between compared 3D objects, we devised a Semantically Aligned Distribution Matching loss where the sorted features in each channel are compared.
+We also employed learnable parameters of the rotation angles to estimate the optimal orientation of synthetic objects.
+The geometric structures and orientations of the synthetic objects are jointly optimized during the dataset distillation.
+Experimental results evaluated on the four widely used benchmark datasets of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] demonstrate that the proposed method outperforms existing distillation methods while maintaining strong cross-architecture generalization.
 Furthermore, we also validated the effectiveness of each component of the proposed method through extensive ablation studies.

### v3

@@
 Abstract
-Large-scale datasets are usually required to train deep neural networks, but it increases the computational complexity hindering the practical applications.
-Recently, dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset to alleviate the computational burden of training while preserving essential task-relevant information.
-However, the dataset distillation for 3D point clouds remains largely unexplored, as the point clouds exhibit fundamentally different characteristics from that of images, making the dataset distillation more challenging.
-In this paper, we propose a distribution matching-based distillation framework for 3D point clouds that jointly optimizes the geometric structures as well as the orientations of the synthetic 3D objects.
-To address the semantic misalignment caused by unordered indexing of points, we introduce a Semantically Aligned Distribution Matching loss computed on the sorted features in each channel.
-Moreover, to address the rotation variation, we jointly learn the optimal rotation angles while updating the synthetic dataset to better align with the original feature distribution.
-Extensive experiments on widely used benchmark datasets demonstrate that the proposed method consistently outperforms existing dataset distillation methods, achieving superior accuracy and strong cross-architecture generalization.
+Large-scale datasets are usually required to train deep neural networks; however, they increase computational complexity, hindering practical applications.
+Recently, dataset distillation for images and texts has attracted considerable attention, as it reduces the original dataset to a small synthetic one to alleviate the computational burden of training while preserving essential task-relevant information.
+However, dataset distillation for 3D point clouds remains largely unexplored, as point clouds exhibit fundamentally different characteristics from those of images, making this task more challenging.
+In this paper, we propose a distribution-matching-based distillation framework for 3D point clouds that jointly optimizes the geometric structures and orientations of synthetic 3D objects.
+To address the semantic misalignment caused by the unordered nature of point clouds, we introduce a Semantically Aligned Distribution Matching (SADM) loss, which is computed on the sorted features within each channel.
+Moreover, to handle rotational variations, we jointly learn optimal rotation angles while updating the synthetic dataset to better align with the original feature distribution.
+Extensive experiments on widely used benchmark datasets demonstrate that the proposed method consistently outperforms existing dataset distillation approaches, achieving higher accuracy and strong cross-architecture generalization.
 1 Introduction
-With the increasing demand for large-scale training datasets, the cost of model retraining from scratch has become a major challenge leading to the research of dataset distillation [21].
+With the increasing demand for large-scale training datasets, the high cost of retraining models from scratch has become a major challenge, motivating research on dataset distillation [dd].
 The objective of dataset distillation is to produce a significantly smaller synthetic dataset from a large original dataset, that preserves the essential task-relevant information contained in the original dataset.
 Hence the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset.
-Existing dataset distillation methods [3, 6, 8, 11, 29, 28, 30, 17, 26, 5, 9] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches.
+Existing dataset distillation methods [mtt1, mtt2, mtt3, att, gm, dm, idm, datadam, m3d, exploit, divers] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches.
 The gradient matching and trajectory matching methods aim to ensure that the synthetic dataset produces similar optimization dynamics to the original dataset.
-To this end, the gradient matching method [29] minimizes the difference between the gradients computed on the synthetic and original datasets.
+To this end, the gradient matching method [gm] minimizes the difference between the gradients computed on the synthetic and original datasets.
 However, it potentially overlooks long-term dependencies in the training process.
-Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on the synthetic dataset follow similar optimization trajectories to those trained on the original dataset.
+Rather than comparing individual gradients, the trajectory matching methods [mtt1, mtt2, mtt3, att] encourage the models trained on the synthetic dataset follow similar optimization trajectories to those trained on the original dataset.
 Note that these methods train the networks while optimizing the synthetic dataset, and significantly increase the computational complexity.
-To alleviate the computational burden in dataset distillation, the distribution matching techniques [28, 30, 17, 26, 5] have been introduced.
-They randomly initialize the networks without training to extract the features, and compare the feature distributions between the original and synthetic datasets.
-While dataset distillation has been extensively studied for structured data such as images [29, 3, 28] and texts [10, 12, 13], its application to 3D point clouds remains almost unexplored.
-PCC [27] simply applied an existing distillation method for images [29] to the 3D point clouds without considering inherent characteristics of 3D point clouds, however it still suffers from high computational complexity.
+To alleviate the computational burden in dataset distillation, the distribution matching techniques [dm, idm, datadam, m3d, exploit] have been introduced.
+The networks are randomly initialized and used without training to extract features, after which the feature distributions of the original and synthetic datasets are compared.
+While dataset distillation has been extensively studied for structured data such as images [gm, mtt1, dm] and texts [textdd1, textdd2, textdd3], its application to 3D point clouds remains almost unexplored.
+PCC [pcc] simply applied an existing distillation method for images [gm] to the 3D point clouds without considering inherent characteristics of 3D point clouds, however it still suffers from high computational complexity.
 Unlike the structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space.
-Therefore, semantically similar regions across different 3D models are often associated with inconsistent orders (indices of points), that makes direct feature comparison incorrect worsening the performance of distribution matching based dataset distillation.
+Therefore, semantically similar regions across different 3D models are often associated with inconsistent orders (i.e., point indices), which makes direct feature comparison incorrect, thereby worsening the performance of distribution-matching-based dataset distillation.
 We refer to this issue as semantic misalignment.
@@
 3D point clouds are often captured or synthesized under arbitrary poses due to the absence of canonical orientations.
-Such rotational difference causes the objects of the same class yield different features from one another, significantly increasing the intra-class variability.
+Such rotational differences cause objects of the same class to produce different features, significantly increasing intra-class variability.
 As a result, it becomes difficult to construct a representative synthetic dataset that faithfully follows the feature distribution of the original dataset.
@@
 The key contributions of this paper are summarized as follows:
-  • To the best of our knowledge, we are the first to propose a distribution matching-based dataset distillation method of 3D point clouds, that jointly optimizes the shapes and orientations of synthetic dataset.
-  • We devised the SADM loss on the sorted features in each channel to preserve the semantic alignment between compared 3D objects.
+  • To the best of our knowledge, we are the first to propose a distribution-matching-based dataset distillation method of 3D point clouds, that jointly optimizes the shapes and orientations of synthetic dataset.
+  • We devised the SADM loss, computed on the sorted features within each channel, to preserve semantic alignment between the compared 3D objects.
   • We validated the proposed method through extensive experiments on the four benchmark datasets widely used for 3D point clouds classification, and showed the superiority of the proposed method over the existing dataset distillation techniques.
@@
 3D point clouds are generally unordered exhibiting irregular characteristics, that makes it difficult to apply the convolution operations in deep neural networks commonly used for images.
-Early approaches [25, 32] convert the point clouds into structured representation of multi-view images or voxel grids to enable the use of conventional deep learning architectures, such as convolutional neural networks (CNNs).
-VoxelNet [32] voxelizes the point clouds, but it causes increased memory consumption and quantization errors.
-To directly process point clouds without conversion to intermediate structured representations, PointNet [14] was introduced as a pioneering work that learns the features directly from unordered raw point clouds.
-However, it suffers from the limitation in capturing local geometric relationships, and hierarchical CNNs have been developed tailored to 3D point clouds.
-PointNet++[15] captures local geometric structures through iterative sampling and grouping strategies.
-On the other hand, PointConv [23] generalizes the convolution operations to irregular point clouds by adaptively weighting the neighboring points based on their spatial distribution.
-Meanwhile, DGCNN [22] constructs dynamic graphs from point clouds to adaptively capture the semantic information.
-Attempts have been also made to apply the transformer [20] to 3D point clouds processing, where the Point Transformer [31] utilizes the attention mechanism to capture the long-range dependencies.
+To process such data, early studies [(multiview)multi, (voxel)voxelnet] converted point clouds into structured representations, such as multi-view images or voxel grids, enabling the use of standard convolutional neural networks (CNN) architectures.
+However, these conversions introduce quantization errors and increase memory overhead.
+PointNet [pointnet] addressed this limitation by directly learning the features from unordered points, which serves as the backbone in our dataset distillation framework.
+Subsequent methods, such as PointNet++ [pointnet2], PointConv [pointconv], and DGCNN [dgcnn], extended PointNet by capturing local geometric relationships through hierarchical architecture.
+Additionally, attempts have been also made to apply the transformer [transformer] to 3D point clouds processing, where Point Transformer [pointtransformer] utilizes the attention mechanism to capture the long-range dependencies.
+We adopt these architectures as evaluation networks to assess the cross-architecture generalization capability of our distilled datasets.
 2.2 Coreset Selection
 Coreset selection is a technique designed to select representative samples from a given dataset, while maintaining the model’s performance even with the sampled data.
-The random selection method [16] randomly chooses a subset of data samples from the whole dataset.
+The random selection method [herding(icarl)] randomly chooses a subset of data samples from the whole dataset.
 It is simple but suffers from the robustness due to the lack of informative criteria for sampling.
-The K-center method [18] selects data points iteratively that maximize the minimum distance to the set of already selected ones, taking into account the data distribution.
-The Herding method [2, 1] iteratively takes data points that minimize the discrepancy between the mean embeddings of the selected subset and the entire dataset in the feature space.
+The K-center method [kcenter] selects data samples iteratively that maximize the minimum distance to the set of already selected ones, taking into account the data distribution.
+The Herding method [herding(e2e), herding(scail)] iteratively takes data points that minimize the discrepancy between the mean embeddings of the selected subset and the entire dataset in the feature space.
+The coreset selection operates within the space of existing samples, and hence lacks the flexibility to synthesize informative patterns.
+This limitation naturally leads to the emergence of dataset distillation, which generates synthetic datasets capturing task-specific patterns beyond the original samples.
 2.3 Dataset Distillation
-Dataset distillation [21] methods can be largely categorized into the gradient matching [29], trajectory matching [3, 6, 8, 11], and distribution matching [28, 30, 17, 26, 5] approaches.
-The gradient matching, first introduced by DC [29], minimizes the difference between the gradients computed from the original and synthetic datasets, respectively, guiding the synthetic dataset to follow the training direction of the original dataset.
-The trajectory matching methods [3, 6, 8, 11], initially proposed by MTT [3], encourage the models trained on the synthetic dataset to follow the optimization trajectories similar to those trained on the original dataset rather than comparing individual gradients.
-ATT [11] automatically adjusts the length of the training trajectories between the synthetic and original datasets, enabling more effective and precise matching.
-The distribution matching, introduced by DM [28], focuses on minimizing the distance between the feature distributions of the original and synthetic datasets.
-DataDAM [17] enhances the distribution matching by aligning the feature maps using attention mechanism, achieving unbiased feature representation with low computational overhead.
-Furthermore, M3D [26] employs the Gaussian kernel function in the distribution matching loss, enabling the alignment across higher-order statistical characteristics of the feature distributions.
-Recently, a gradient matching-based point clouds distillation method [27] has been introduced, however it simply applied the existing image-based method without considering the unique characteristics of 3D point clouds.
+Dataset distillation [dd] methods can be largely categorized into the gradient matching [gm], trajectory matching [mtt1, mtt2, mtt3, att], and distribution matching [dm, idm, datadam, m3d, exploit] approaches.
+The gradient matching, first introduced by DC [gm], minimizes the difference between the gradients computed from the original and synthetic datasets, respectively, guiding the synthetic dataset to follow the training direction of the original dataset.
+The trajectory matching methods [mtt1, mtt2, mtt3, att], initially proposed by MTT [mtt1], encourage the models trained on the synthetic dataset to follow the optimization trajectories similar to those trained on the original dataset rather than comparing individual gradients.
+ATT [att] automatically adjusts the length of the training trajectories between the synthetic and original datasets, enabling more effective and precise matching.
+The distribution matching, introduced by DM [dm], focuses on minimizing the distance between the feature distributions of the original and synthetic datasets.
+DataDAM [datadam] enhances the distribution matching by aligning the feature maps using attention mechanism, achieving unbiased feature representation with low computational overhead.
+Furthermore, M3D [m3d] employs the Gaussian kernel function in the distribution matching loss, enabling the alignment across higher-order statistical characteristics of the feature distributions.
+Recently, a gradient matching-based point clouds distillation method [pcc] has been introduced, however it simply applied the existing image-based method without considering the unique characteristics of 3D point clouds.
 3 Methodology
@@
 Problem Definition.
-The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}_{i=1}^{|\mathcal{T}|}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}_{i=1}^{|\mathcal{S}|}$, where $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, such that a model trained on ${\boldsymbol{\mathcal{S}}}$ achieves close performance to that trained on ${\boldsymbol{\mathcal{T}}}$.
+The objective of dataset distillation is to compress the task-relevant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}_{i=1}^{|\mathcal{T}|}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}_{i=1}^{|\mathcal{S}|}$, where $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, such that a model trained on $\boldsymbol{\mathcal{S}}$} achieves close performance to that trained on $\boldsymbol{\mathcal{T}}$}.
 Given a 3D point cloud sample $\bf p$}} following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset $\boldsymbol{\mathcal{S}}$$^{\star}$ can be obtained via
-${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}\limits_{{\boldsymbol{\mathcal{S}}}}\mathbb{E}_{{{\bf p}}}\left[||{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{T}}}}}({{\bf p}}),l)-{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{S}}}}}({{\bf p}}),l)||^{2}\right],$ (1)
+[…]
 where $\mathcal{L}$} denotes a task-specific loss function, such as the cross-entropy loss, and $\phi_{{\boldsymbol{\mathcal{T}}}}$ and $\phi_{{\boldsymbol{\mathcal{S}}}}$ represent the models to estimate the class label which are trained on $\boldsymbol{\mathcal{T}}$} and $\boldsymbol{\mathcal{S}}$}, respectively.
-Distribution Matching. Distribution matching (DM) strategy focuses on aligning the feature distributions derived from the original and synthetic datasets, respectively, via
+Distribution Matching. The distribution matching (DM) strategy focuses on aligning the feature distributions derived from the original and synthetic datasets, respectively, via
 ${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}_{{\boldsymbol{\mathcal{S}}}}D(\phi({\boldsymbol{\mathcal{T}}}),\phi({\boldsymbol{\mathcal{S}}})),$ (2)
 where $\phi$ is the feature extractor and $D$ denotes a distance function.
-We employ a randomly initialized, untrained network, which has been demonstrated to sufficiently capture the structural information for distribution alignment [28].
-Also, the Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used as $D$, given by
+We employ a randomly initialized, untrained network, which has been demonstrated to sufficiently capture the structural information for distribution alignment [dm].
+Also, the Maximum Mean Discrepancy (MMD) [mmd] loss $\mathcal{L}_{\text{MMD}}$ is often used as $D$, given by
 $\mathcal{L}_{\text{MMD}}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}}},{\boldsymbol{\mathcal{S}}})-2{K}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{S}}}),$ (3)
@@
 Specifically, given a 3D point cloud object $\bf p$}}, we extract the features using $\phi$ via
-$\phi({{\bf p}})=\left[{\bf f}_{1},{\bf f}_{2},\dots,{\bf f}_{C}\right],\quad{\bf f}_{i}\in\mathbb{R}^{N},$ (5)
+$\phi({{\bf p}})=\begin{bmatrix}f_{1,1}&f_{2,1}&\cdots&f_{C,1}\\ f_{1,2}&f_{2,2}&\cdots&f_{C,2}\\ \vdots&\vdots&\ddots&\vdots\\ f_{1,N}&f_{2,N}&\cdots&f_{C,N}\end{bmatrix}=\left[{\bf f}_{1},{\bf f}_{2},\dots,{\bf f}_{C}\right],\quad{\bf f}_{i}\in\mathbb{R}^{N},$ (5)
 where $N$ denotes the number of points in $\bf p$}}, $C$ is the number of feature channels, and ${\bf f}_{i}=\left\{f_{i,1},f_{i,2},...,f_{i,N}\right\}$ represents the feature vector of the $i$-th channel where $f_{i,j}$ is the feature value of the $j$-th point.
-We then perform channel-wise sorting to the extracted features of $f_{i,j}$’s in the descending order of their sizes, and obtain the sorted feature vector $\tilde{{\bf f}}_{i}=\left\{\tilde{f}_{i,1},\tilde{f}_{i,2},...,\tilde{f}_{i,N}\right\}$ such that $\tilde{f}_{i,j}\geq\tilde{f}_{i,j+1}$.
+We then perform channel-wise sorting to the extracted features of $f_{i,j}$’s in the descending order of their values, and obtain the sorted feature vector $\tilde{{\bf f}}_{i}=\left\{\tilde{f}_{i,1},\tilde{f}_{i,2},...,\tilde{f}_{i,N}\right\}$ such that $\tilde{f}_{i,j}\geq\tilde{f}_{i,j+1}$.
 Then we have the set of the sorted feature vectors as
@@
 To measure the discrepancy between the sorted features, we redefine the Gaussian kernel as
-$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\exp\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
+[…]
 Then we devise the loss $\mathcal{L}_{\alpha}$ from $\mathcal{L}_{\text{MMD}}$ (3), given by
@@
 $\mathcal{L}_{\beta}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})=\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}})+\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}})-2\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (9)
-where $\tilde{K}_{\text{top}}(\cdot,\cdot)$ denotes the Gaussian kernel computed with only the most largest feature in each channel.
+where $\tilde{K}_{\text{top}}(\cdot,\cdot)$ denotes the Gaussian kernel computed with only the largest feature in each channel.
 The SADM loss is then formulated as a weighted sum of the two losses, given by
@@
 where $\lambda_{1},\lambda_{2}$ are the weighting parameters.
-By using $\mathcal{L}_{\text{SADM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we facilitate reliable matching of feature distributions considering sematic structures of 3D point clouds, thereby improving the performance dataset distillation.
+By using $\mathcal{L}_{\text{SADM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we facilitate reliable matching of feature distributions considering semantic structures of 3D point clouds, thereby improving the performance dataset distillation.
 3.3 Estimation of Optimal Rotations
@@
 These angles are treated as learnable parameters, allowing the orientation of each synthetic 3D object in $\boldsymbol{\mathcal{S}}$} to be adaptively adjusted during dataset distillation.
-Therefore, instead of minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (10), we minimize $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset whose 3D objects are rotated toward the estimated optimal orientations.
+Therefore, instead of minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (10), we minimize $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})})$, where $\mathcal{R}_{\boldsymbol{\theta}}$ denotes the rotation operator according to the rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$.
+The overall objective is formally defined as
+$\left\{{\boldsymbol{\mathcal{S}}}^{\star},{\boldsymbol{\theta}}^{\star}\right\}=\mathop{\arg\min}_{\left\{{\boldsymbol{\mathcal{S}}},{\boldsymbol{\theta}}\right\}}\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})).$ (11)
 Note that we optimize the synthetic 3D objects as well as their rotation parameters simultaneously, during the dataset distillation process.
 Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset.
-We then form $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the objects in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to zero.
+We then form $\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})$ by rotating the objects in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to zero.
 For each class, a mini-batch is sampled from $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per class.
-The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with the batch size determined by the number of point cloud objects per class (PPC).
-The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and their rotation parameters by minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$.
+The corresponding synthetic mini-batch is sampled from $\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})$, with the batch size determined by the number of point cloud objects per class (PPC).
+The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and their rotation parameters by minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})})$.
 This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while aligning their orientations more effectively.
@@
 Proposition 1.
-Jointly optimizing the synthetic dataset $\boldsymbol{\mathcal{S}}$ and the rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ guarantees a lower or equal loss to that of optimizing $\boldsymbol{\mathcal{S}}$ alone.
-$\min_{\boldsymbol{\mathcal{S}},\boldsymbol{\theta}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}}))\leq\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (11)
+Jointly optimizing the synthetic dataset $\boldsymbol{\mathcal{S}}$ and the rotation parameters $\boldsymbol{\theta}$ guarantees a lower or equal loss to that of optimizing $\boldsymbol{\mathcal{S}}$ alone.
+$\min_{\left\{\boldsymbol{\mathcal{S}},\boldsymbol{\theta}\right\}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}}))\leq\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (12)
 where $\mathcal{R}_{\boldsymbol{\theta}}$ denotes the rotation operator according to $\boldsymbol{\theta}$.
@@
 The proof is provided in Appendix A.
-Allowing optimization over both $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ enlarges the feasible set compared to optimizing only $\boldsymbol{\mathcal{S}}$.
-The feasible set for the joint optimization includes the feasible set of the original optimization as a subset, since optimizing only $\boldsymbol{\mathcal{S}}$ is equivalent to the special case of the joint optimization with the fixed $(\theta_{x},\theta_{y},\theta_{z})=(0,0,0)$.
-Consequently, the minimum loss achieved by jointly optimizing over $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ must be less than or equal to the minimum loss achieved by optimizing only $\boldsymbol{\mathcal{S}}$.
+Jointly optimizing over $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ enlarges the feasible set, as the original optimization over $\boldsymbol{\mathcal{S}}$ alone is a special case of the joint optimization with fixed $(\theta_{x},\theta_{y},\theta_{z})=(0,0,0)$.
 $\square$
 Table 1: Comparison of quantitative performance.
-DC, DM, and the proposed method were initialized with random selection [16] for fair comparison.
+DC, DM, and the proposed method were initialized with random selection [herding(icarl)] for fair comparison.
+‘Whole’ refers to the classification accuracy obtained by training the network on the entire original dataset without any distillation.
+‘Ratio’ represents the percentage of the size of the distilled dataset compared to that of the original dataset.
 The best and the second best scores are highlighted in bold and underlined, respectively.
-Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
+Datasets ModelNet10 [modelnet40] ModelNet40 [modelnet40] ShapeNet [shapenet] ScanObjectNN [scanobjectnn]
 PPC 1 3 10 1 3 10 1 3 10 1 3 10
@@
 Whole 91.41 87.84 82.49 63.84
-Random 35.5±$\pm 4.7$ 75.2±$\pm 1.7$ 85.3±$\pm 1.1$ 34.6±$\pm 1.8$ 60.0±$\pm 1.32$ 74.1±$\pm 0.4$ 34.0±$\pm 3.0$ 54.8±$\pm 1.5$ 63.1±$\pm 1.0$ 13.9±$\pm 1.4$ 20.4±$\pm 1.3$ 34.8±$\pm 1.1$
-Herding 40.1 ±$\pm 5.2$ 78.0 ±$\pm 1.3$ 86.9±$\pm 0.6$ 54.4 ±$\pm 2.0$ 68.5 ±$\pm 1.0$ 78.8 ±$\pm 0.4$ 49.5 ±$\pm 2.3$ 59.8 ±$\pm 0.9$ 66.9 ±$\pm 0.5$ 15.7 ±$\pm 1.7$ 27.7 ±$\pm 1.4$ 38.7 ±$\pm 1.7$
-K-center 40.1 ±$\pm 5.2$ 77.6±$\pm 1.9$ 83.2±$\pm 1.4$ 54.4±$\pm 2.0$ 63.0 ±$\pm 2.7$ 65.3 ±$\pm 1.1$ 49.5 ±$\pm 2.3$ 51.4 ±$\pm 1.7$ 47.8 ±$\pm 0.7$ 15.7 ±$\pm 1.7$ 19.8 ±$\pm 0.8$ 24.0 ±$\pm 1.0$
-DM 31.9±$\pm 4.3$ 77.6±$\pm 1.6$ 86.1±$\pm 0.8$ 32.3±$\pm 4.8$ 62.0±$\pm 1.9$ 75.2±$\pm 0.5$ 27.8±$\pm 4.0$ 56.1±$\pm 1.2$ 64.3±$\pm 0.7$ 14.9±$\pm 2.0$ 23.3±$\pm 1.5$ 35.6±$\pm 1.4$
-DC 43.9±$\pm 5.5$ 75.0±$\pm 2.4$ 86.1±$\pm 1.1$ 52.0±$\pm 1.7$ 66.6±$\pm 1.3$ 75.6±$\pm 0.5$ 49.8±$\pm 1.3$ 60.1±$\pm 1.4$ 64.6±$\pm 0.7$ 18.6±$\pm\textbf{1.6}$ 24.3±$\pm 2.5$ 35.3±$\pm 1.2$
-Ours 44.7 ±$\pm\textbf{6.1}$ 84.4 ±$\pm\textbf{1.2}$ 87.8 ±$\pm\textbf{1.0}$ 55.8 ±$\pm\textbf{1.5}$ 72.1 ±$\pm\textbf{0.8}$ 80.1 ±$\pm\textbf{0.4}$ 50.2 ±$\pm\textbf{1.4}$ 63.7 ±$\pm\textbf{0.7}$ 68.4 ±$\pm\textbf{0.5}$ 17.4 ±$\pm 1.5$ 31.6 ±$\pm\textbf{1.0}$ 43.9 ±$\pm\textbf{1.9}$
-Table 2: Comparison of cross-architecture generalization performance evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
-Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
+[…]
+[…]
+[…]
+[…]
+[…]
+Ours 44.7 $\pm\textbf{6.1}$ 84.4 $\pm\textbf{1.2}$ 87.8 $\pm\textbf{1.0}$ 55.8 $\pm\textbf{1.5}$ 72.1 $\pm\textbf{0.8}$ 80.1 $\pm\textbf{0.4}$ 50.2 $\pm\textbf{1.4}$ 63.7 $\pm\textbf{0.7}$ 68.4 $\pm\textbf{0.5}$ 17.4 $\pm 1.5$ 31.6 $\pm\textbf{1.0}$ 43.9 $\pm\textbf{1.9}$
+Table 2: Comparison of cross-architecture generalization performance evaluated on PointNet++ [pointnet2], DGCNN [dgcnn], PointConv [pointconv], and PT [pointtransformer], respectively.
+Datasets ModelNet10 [modelnet40] ModelNet40 [modelnet40] ShapeNet [shapenet] ScanObjectNN [scanobjectnn]
 Method DM DC Ours DM DC Ours DM DC Ours DM DC Ours
-PointNet++ 35.0±$\pm 8.3$ 51.5±$\pm 7.8$ 82.9±$\pm\textbf{1.4}$ 14.1±$\pm 2.7$ 46.9±$\pm 6.0$ 68.5±$\pm\textbf{0.5}$ 20.3±$\pm 3.7$ $\pm 3.4$ 53.2±$\pm\textbf{1.5}$ 15.5±$\pm 3.6$ 11.6±$\pm 2.2$ 22.0±$\pm\textbf{3.9}$
-DGCNN 56.6±$\pm 5.3$ 62.6±$\pm 2.2$ 79.0±$\pm\textbf{2.2}$ 34.8±$\pm 3.5$ 52.3±$\pm 2.7$ 66.9±$\pm\textbf{1.1}$ 13.8±$\pm 2.2$ 38.0±$\pm 2.7$ 52.8±$\pm\textbf{1.6}$ 17.7±$\pm 2.4$ 12.5±$\pm 1.3$ 19.0±$\pm\textbf{1.5}$
-PointConv 33.4±$\pm 6.2$ 40.0±$\pm 10.5$ 56.6±$\pm\textbf{4.9}$ 20.0±$\pm 5.7$ 37.9±$\pm 3.9$ 51.3±$\pm\textbf{6.2}$ 17.4±$\pm 3.0$ 34.0±$\pm 4.8$ 47.4±$\pm\textbf{2.3}$ 15.2±$\pm 1.7$ 14.2±$\pm 1.8$ 16.8±$\pm\textbf{1.5}$
-PT 48.5±$\pm 6.6$ $\pm 1.6$ 77.6±$\pm\textbf{1.6}$ 26.7±$\pm 5.0$ 47.5±$\pm 2.7$ 61.6±$\pm\textbf{1.1}$ 34.2±$\pm 5.3$ 42.6±$\pm 2.5$ 55.5±$\pm\textbf{0.5}$ 17.1±$\pm 0.8$ 14.9±$\pm 0.5$ 21.9±$\pm\textbf{2.7}$
+[…]
+[…]
+[…]
+[…]
 4 Experimental Results
 4.1 Experimental Setup
-The proposed method was evaluated on the ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets.
+The proposed method was evaluated on the ModelNet10 [modelnet40], ModelNet40 [modelnet40], ShapeNet [shapenet], and ScanObjectNN [scanobjectnn] datasets.
 The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively.
 The ScanObjectNN consists of real-world scanned objects from 15 classes.
-Following previous methods [29, 28], we measure the classification accuracy trained on the distilled synthetic dataset.
+Following previous methods [gm, dm], we measure the classification accuracy trained on the distilled synthetic dataset.
 We ensure the fairness by training the network 10 times and report the mean and standard deviation of accuracy.
-We evaluate the performance across different PPC values of 1, 3, and 10, respectively.
-Each point cloud object contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15].
-The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iteration by training the PointNet on the synthetic dataset and testing it on the original test dataset.
+We evaluate the performance across different PPC values of 1, 3, and 10.
+Each point cloud object contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [pointnet, pointnet2].
+The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iterations by training the PointNet on the synthetic dataset and testing it on the original test dataset.
 4.2 Performance Comparison
 4.2.1 Dataset Distillation
-We compare the performance of the proposed method with that of 1) the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18], 2) two existing image dataset distillation methods of DC [29] and DM [28], and 3) the only existing 3D point clouds dataset distillation method of PCC [27].
+We compare the performance of the proposed method with that of 1) the three most representative coreset selection methods: random selection [herding(icarl)], Herding [herding(e2e), herding(scail)], and K-Center [kcenter], 2) two existing image dataset distillation methods of DC [gm] and DM [dm].
 Table 1 compares the quantitative performance of the proposed method with that of the existing methods.
 DM and DC were initialized with random selection.
-The synthetic datasets were optimized using PointNet [14], and classification performance was evaluated on PointNet to demonstrate their validity.
+The synthetic datasets were optimized using PointNet [pointnet], and classification performance was evaluated on PointNet to demonstrate their validity.
 Among the coreset selection methods, Herding achieves the best overall performance.
@@
 As PPC increases, Herding consistently outperforms other coreset selection methods, showing its effectiveness.
-However, since the coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions.
+However, because coreset selection methods rely solely on sample selection without further optimization, they struggle to capture complex feature distributions.
 The features used in DM are extracted from the layer before the classifier, and they discard substantial information through the global max pooling process of PointNet.
@@
 As a result, it consistently underperforms compared to the proposed method.
-The proposed method outperforms both the coreset selection and existing distillation techniques.
+The proposed method outperforms both coreset selection and existing distillation techniques.
 When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement.
 As the PPC increases, the proposed method more effectively matches the feature distributions significantly improving the performance.
+Table 3: Comparison of cross-dataset generalization performance at the PPC value of 3, where the models are trained on a subset of ModelNet40 and evaluated on a subset of ShapeNet, and vice versa.
+Setting DC DM PCC Ours
+SN $\rightarrow$ MN40 54.43 49.66 57.43 61.54
+MN40 $\rightarrow$ SN 53.75 47.35 49.42 54.40
+Table 5: Comparison of training times (in hours) required for dataset distillation at the PPC value of 3.
+Method MN10 MN40 SN SONN
+DC 1.52 5.43 7.57 2.19
+DM 0.04 0.11 0.15 0.05
+Ours 0.08 0.26 0.35 0.11
 Table 4: Performance comparison across different initialization strategies.
-The accuracies of three PPC values of 1, 3, and 10 are averaged.
+The accuracies for PPC values of 1, 3, and 10 are averaged.
 MN10: ModelNet10.
@@
 Ours 72.93 69.67 62.25 31.08
-† DC [29] with Herding initialization corresponds to PCC [27]
-Table 4: Comparison of training time (in hours) at the PPC value of 3.
-Method MN10 MN40 SN SONN
-DC 1.52 5.43 7.57 2.19
-DM 0.04 0.11 0.15 0.05
-Ours 0.08 0.26 0.35 0.11
-Table 5: Ablation study of the proposed SADM loss.
-Experiments were performed at the PPC value of 3.
-“Random” denotes the results of the random selection [16], and SA is the proposed semantic alignment.
-MN10 MN40 SN SONN
-Random 75.17 59.96 54.84 20.42
-w/o SA 75.01 59.96 53.91 20.20
-w/ SA 84.42 72.08 63.74 31.84
+† DC [gm] with Herding initialization corresponds to PCC [pcc]
 4.2.2 Cross-Architecture Generalization
-We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31], respectively.
-Unlike PointNet, the architectures used for generalization performance comparison follow hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from that of PointNet.
+We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [pointnet] and evaluated on different backbone networks including PointNet++[pointnet2], DGCNN[dgcnn], PointConv [pointconv], and PT [pointtransformer], respectively.
+Unlike PointNet, the architectures used for generalization performance comparison follow hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from those of PointNet.
 This difference makes it essential for the distilled dataset to retain rich and transferable features that generalize well toward different architectures.
@@
 Wherease DC suffers from the overfitting and DM loses critical information, the proposed method yields higher adaptability showing strong generalization performance across various backbone networks.
-Figure 3: Comparison of synthetic datasets distilled from the ModelNet40 [24] dataset.
-Point clouds were colorized according to the y-coordinates.
-From top to bottom, the orginal dataset (first), initialized point cloud models (second), and the distilled synthetic datasets by using DC [29] (third), DM [28] (fourth), and the proposed method (fifth).
+4.2.3 Cross-Dataset Generalization
+To further evaluate the generalization capability of the distilled datasets across different domains, we provide an experimental result of cross-dataset experiment in Table 4.
+Specifically, we constructed the subsets of ShapeNet and ModelNet40, respectively, by selecting the 16 object classes shared between them.
+Then we performed the experiments in which the dataset was distilled on the subset of ShapeNet and tested on ModelNet40 (SN $\rightarrow$ MN40), and vice versa (MN40 $\rightarrow$ SN).
+As shown in Table 4, the proposed method consistently achieves the best performance in both settings of the cross-dataset evaluation, indicating that our distilled datasets effectively capture geometric patterns that are useful regardless of the datasets.
 4.2.4 Initialization Strategies
-Table 5 presents the classification performance on PointNet for DC [29], DM [28], and PCC [27] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18].
-Note that the results of PCC are obtained by our implementation, where the PCC is equivalent to performing DC with Herding initialization.
+Table 4 presents the classification performance of DC [gm], DM [dm], and PCC [pcc] under different initialization strategies, including uniform noise, random selection [herding(icarl)], Herding [herding(e2e), herding(scail)], and K-Center [kcenter].
+Note that the results for PCC are obtained using our implementation, where PCC is equivalent to performing DC with Herding initialization.
 When initialized with noise, DC fails to effectively optimize the synthetic dataset because the network has far more parameters than typical networks used for image dataset distillation, making it difficult to align the gradients.
@@
 In contrast, the structured initializations such as random selection, Herding, or K-Center improve the performance, where the Herding performs the best by selecting the most representative samples.
-Unlike DC, DM directly matches the feature representations rather than relying on the gradients and hence less sensitive to the initialization strategies.
-However, due to the information loss, DM consistently underperforms DC even with the structured initializations.
-In contrast, the proposed method achieves favorable performance, even when initialized from uniform noise.
-Furthermore, the proposed method significantly outperforms other approaches under the heuristic initialization strategy, such as random selection.
-This demonstrates that the proposed method is robust with diverse initialization schemes and effectively resolves the instability issues inherent in the existing dataset distillation methods.
+Unlike DC, DM directly matches feature representations rather than relying on gradients, making it less sensitive to initialization strategies.
+However, due to information loss, DM consistently underperforms DC, even when structured initializations are used.
+In contrast, the proposed method achieves favorable performance even when initialized from uniform noise.
+Furthermore, it significantly outperforms other approaches under heuristic initialization strategies, such as random selection.
+These results demonstrate that the proposed method is robust across diverse initialization schemes and effectively addresses the instability issues inherent in existing dataset distillation methods.
 4.2.5 Training Time
 Table 4 compares the training times between the proposed method and the existing methods.
-DM [28] achieves the fastest training time as it uses only a small subset of the overall features, however it often ignores many useful features leading to performance degradation.
-In contrast, DC [29] incurs a significantly higher computational cost as it aligns the gradients across all the parameters.
-Notably, for the largest dataset of ShapeNet [4], DC requires over 7 hours for training 50 times slower that DM that completes the training in just 9 minutes (0.15 hours).
+DM [dm] achieves the fastest training time as it uses only a small subset of the overall features, however it often ignores many useful features leading to performance degradation.
+In contrast, DC [gm] incurs a significantly higher computational cost as it aligns the gradients across all the parameters.
+Notably, for the largest dataset of ShapeNet [shapenet], DC requires over 7 hours for training 50 times slower that DM that completes the training in just 9 minutes (0.15 hours).
 Since the proposed method is based on the distribution matching approach, its training time is marginally increased compared to DM, yet achieves significant improvement in dataset distillation performance.
+To further justify the computational overhead, we conducted an experiment where DM was trained with its default setting, whereas the proposed method was trained during the same training time as DM.
+As shown in Table 7, even under the same training time condition, our method consistently outperforms DM.
+Table 6: Performance comparison between DM and the proposed method at the PPC value of 3, where the proposed method was trained during the same time as DM.
+Method MN10 MN40 SN SONN
+DM 77.58 62.04 56.14 23.33
+Ours 83.64 70.56 63.34 33.06
+Table 7: Ablation study of the proposed semantic alignment (SA) method. Experiments were performed at the PPC value of 3.
+Method MN10 MN40 SN SONN
+Random 75.17 59.96 54.84 20.42
+w/o SA 75.01 59.96 53.91 20.20
+w/ SA 84.42 72.08 63.74 31.84
 4.2.6 Qualitative Comparison of Synthetic Datasets
@@
 The first row shows the original datasets and the second row shows initialized point cloud objects.
-DC [29] fails to deviate significantly from the initialized objects and causes noise.
-Similarly, DM [28] also maintains the original shape of the initialized objects while shifting certain points only, that hinders to capture meaningful structural changes.
+DC [gm] fails to deviate significantly from the initialized objects and causes noise.
+Similarly, DM [dm] also maintains the original shape of the initialized objects while shifting certain points only, that hinders to capture meaningful structural changes.
 In contrast, as shown in the bottom row, the proposed method successfully preserves the overall semantic structures of 3D objects in each class, while selectively learning essential features.
@@
 Also, in the guitar class, the proposed method jointly optimizes the shape and orientation of the synthesized 3D objects.
+Figure 3: Comparison of synthetic datasets distilled from the ModelNet40 [modelnet40] dataset.
+Point clouds were colorized according to the y-coordinates.
+From top to bottom, the orginal dataset (first), initialized point cloud models (second), and the distilled synthetic datasets by using DC [gm] (third), DM [dm] (fourth), and the proposed method (fifth).
 4.3 Ablation Study
 4.3.1 Effect of SADM Loss
-Table 5 presents the ablation study on the impact of the proposed SADM loss on the classification performance using PointNet [14].
-Without preserving the semantic alignment, the unordered nature of point clouds restricts effective feature matching, resulting in nearly identical performance to that of the random selection [16].
-That means the model has not been properly optimized.
-Otherwise, when the semantic alignment is applied, accuracy improves across all datasets, as the aligned featured provide consistent and structured representation that facilitates more effective feature matching during the optimization.
-This result confirms that preserving the inherent semantic alignment via feature sorting facilitates the network to exploit well-aligned corresponding features and significantly improves the classification performance.
+Table 7 presents an ablation study on the impact of the proposed SADM loss on classification performance using PointNet [pointnet].
+Without preserving semantic alignment, the unordered nature of point clouds restricts effective feature matching, resulting in performance nearly identical to that of random selection [herding(icarl)].
+This indicates that the model has not been properly optimized.
+In contrast, when semantic alignment is applied, accuracy improves across all datasets, as the aligned features provide a consistent and structured representation that facilitates more effective feature matching during optimization.
+These results confirm that preserving inherent semantic alignment via feature sorting enables the network to exploit well-aligned corresponding features, significantly enhancing classification performance.
 4.3.2 Effect of Optimal Rotation Estimation
+Table 9 presents the ablation study to evaluate the impact of joint optimization of synthetic dataset and rotation angles.
+We categorized the test datsets into the aligned, mixed, and rotated groups to analyze the effectiveness of the proposed rotation estimation.
+The aligned group consists of the datasets where the objects maintain consistent orientations, including ModelNet10 and ShapeNet [shapenet].
+The mixed group includes ModelNet40 [modelnet40] where only certain classes exhibit rotation variations.
+The rotated group is ScanObjectNN [scanobjectnn] where the objects exhibit arbitrary rotations across all classes.
+The baseline (first row) shows the performance without optimization which is identical to that of the random selection.
+In the second row, we present the results where only the orientation of the synthetic data is optimized using the proposed SADM loss, without updating the synthetic dataset.
+This leads to performance improvements across all the groups compared to random selection.
+The improvement becomes more significant as the classes with rotation variation become more dominant.
+Optimizing the synthetic dataset further boosts the performance, with the best results achieved when both the synthetic data and rotation angles are optimized simultaneously.
+In particular, the accuracy on the rotated dataset improves significantly, demonstrating that the proposed joint optimization enables the model to be resilient to rotation variations.
+4.3.3 Effect of Point Sorting
+We compared the performance of the proposed SADM against two representative point sorting schemes.
+The axis-aligned sorting method simply orders the points in the ascending order of their $z$-coordinates, without considering any structural or semantic relationships.
+The Z-order sorting method [morton] assigns an index to each point based on its 3D coordinates simultaneously, ensuring that points close in space receive nearby indices.
+As shown in Table 9, the existing sorting methods yield only marginal improvements over the baseline that does not use point sorting.
+Moreover, they fail to provide consistent semantic alignment across different samples.
+In contrast, the proposed SADM significantly outperforms these methods by achieving semantically consistent matching, thereby enhancing the effectiveness of dataset distillation for 3D point clouds.
 Table 8: Ablation study of the proposed optimal rotation estimation. Experiments were performed at the PPC value of 3.
@@
 74.35 72.08 31.84
-Table 6 presents the ablation study to evaluate the impact of joint optimization of synthetic dataset and rotation angles.
-We categorized the test datsets into the aligned, mixed, and rotated groups to analyze the effectiveness of the proposed rotation estimation.
-The aligned group consists of the datasets where the objects maintain consistent orientations, including ModelNet10 and ShapeNet [4].
-The mixed group includes ModelNet40 [24] where only certain classes exhibit rotation variations.
-The rotated group is ScanObjectNN [19] where the objects exhibit arbitrary rotations across all classes.
-The baseline (first row) shows the performance without optimization which is identical to that of the random selection.
-In the second row, we present the results where only the orientation of the synthetic data is optimized using the proposed SADM loss, without updating the synthetic dataset.
-This leads to performance improvements across all the groups compared to random selection.
-The improvement becomes more significant as the classes with rotation variation become more dominant.
-Optimizing the synthetic dataset further boosts the performance, with the best results achieved when both the synthetic data and rotation angles are optimized simultaneously.
-In particular, the accuracy on the rotated dataset improves significantly, demonstrating that the proposed joint optimization enables the model to be resilient to rotation variations.
+Table 9: Performance comparison of different point sorting schemes at the PPC value of 3. For fair comparison, the rotation parameter optimization was not applied.
+Method MN10 MN40 SN SONN
+Unsorted 75.21 59.96 53.91 20.20
+Axis-Aligned 76.44 60.85 53.97 21.80
+Z-order [morton] 76.55 61.47 55.17 21.76
+SADM 83.17 67.52 60.71 27.66
 4.4 Limitation
 While the proposed SADM loss enhances the semantic consistency by sorting the features, it does not completely align semantically meaningful regions across different point cloud objects.
-Estimating optimal rotations also improves the complexity, particularly when the datasets are already well-aligned to the canonical axes.
+Estimating optimal rotations also increases the complexity, particularly when the datasets are already well-aligned to the canonical axes.
 5 Conclusion
-We proposed a semantically-aligned and orientation-aware dataset distillation framework for 3D point clouds.
-To alleviate the inconsistency of point ordering between compared 3D objects, we devised a Semantically Aligned Distribution Matching loss where the sorted features in each channel are compared.
-We also employed learnable parameters of the rotation angles to estimate the optimal orientation of synthetic objects.
-The geometric structures and orientations of the synthetic objects are jointly optimized during the dataset distillation.
-Experimental results evaluated on the four widely used benchmark datasets of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] demonstrate that the proposed method outperforms existing distillation methods while maintaining strong cross-architecture generalization.
-Furthermore, we also validated the effectiveness of each component of the proposed method through extensive ablation studies.
+We proposed a semantically aligned and orientation-aware dataset distillation framework for 3D point clouds.
+To address the inconsistency in point ordering between compared 3D objects, we devised a Semantically Aligned Distribution Matching (SADM) loss that compares sorted features within each channel.
+Additionally, we introduced learnable rotation angle parameters to estimate the optimal orientation of synthetic objects.
+The geometric structures and orientations of the synthetic objects are jointly optimized during dataset distillation.
+Experimental results on four widely used benchmark datasets—ModelNet10 [modelnet40], ModelNet40 [modelnet40], ShapeNet [shapenet], and ScanObjectNN [scanobjectnn]—demonstrate that the proposed method outperforms existing distillation approaches while maintaining strong cross-architecture generalization.
+Furthermore, the effectiveness of each component is validated through extensive ablation studies.
+Acknowledgement
+This work was supported in part by the National Research Foundation of Korea (NRF) Grant funded by Korea Government [Ministry of Science and ICT (MSIT)] under Grant RS-2024-00392536; in part by the Institute of Information and Communications Technology Planning and Evaluation (IITP) Grant funded by Korea Government (MSIT) (Leading Generative AI Human Resources Development) under Grant IITP-2025-RS-2024-00360227; in part by the Artificial Intelligence Gradate School Program, Ulsan National Institute of Science of Technology, under Grant RS-2020-II201336; and in part by the Artificial Intelligence Innovation Hub under Grant RS-2021-II212068.
+Appendix
+Appendix A Proof of Proposition 1
+Proposition 1.
+Jointly optimizing the synthetic dataset $\boldsymbol{\mathcal{S}}$ and the rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ guarantees a lower or equal loss to that of optimizing $\boldsymbol{\mathcal{S}}$ alone.
+$\min_{\left\{\boldsymbol{\mathcal{S}},\boldsymbol{\theta}\right\}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}}))\leq\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (13)
+where $\mathcal{R}_{\boldsymbol{\theta}}$ denotes the rotation operator according to $\boldsymbol{\theta}$.
+Proof. Let us consider two optimization objectives. The first is the baseline optimization, which optimizes only the synthetic dataset without applying any rotation
+$L:=\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}\left(\boldsymbol{\mathcal{T}},\,\boldsymbol{\mathcal{S}}\right).$ (14)
+The second is the proposed joint optimization over both the synthetic dataset $\boldsymbol{\mathcal{S}}$ and the rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$, given
+$L^{\star}:=\min_{\left\{\boldsymbol{\mathcal{S}},\,\boldsymbol{\theta}\right\}}\mathcal{L}_{\text{SADM}}\left(\boldsymbol{\mathcal{T}},\,\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})\right).$ (15)
+Consider a special case of the proposed joint optimization framework where no rotation is applied, i.e., the rotation parameters are fixed to $\boldsymbol{\theta}=(0,0,0)$.
+In such a case, $\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})=\boldsymbol{\mathcal{S}}$, and the baseline optimization can be regarded as a special case of the proposed joint optimization.
+This implies that the feasible set of the baseline optimization is a subset of that of the joint optimization such that
+$\left\{(\boldsymbol{\mathcal{S}},\,\boldsymbol{\theta})\,\middle|\,\boldsymbol{\theta}=(0,0,0)\right\}\subseteq\left\{(\boldsymbol{\mathcal{S}},\,\boldsymbol{\theta})\right\}.$ (16)
+The joint optimization is performed over a larger feasible set than that of the baseline optimization, and therefore its optimal solution must not be higher than that of the baseline. Therefore,
+$L^{\star}\leq L.$ (17)
+Appendix B Expeimental Details
+Datasets. Table 10 summarizes the number of training and testing samples for ModelNet10 [modelnet40], ModelNet40 [modelnet40], ShapeNet [shapenet], and ScanObjectNN [scanobjectnn].
+  • ModelNet10 consists of 10 categories of aligned 3D CAD models. There is no rotational variation.
+  • ModelNet40 includes 40 categories from the same source to ModelNet10, with partially misaligned instances introducing moderate rotational variation.
+  • ShapeNet contains 55 categories of manually aligned 3D models without rotational variation.
+  • ScanObjectNN consists of 15 categories from real-world RGB-D scans with background clutter, occlusion, and sensor noise.
+Objects are unaligned, and rotation variation is inherent.
+We use the PB_T50_RS split, which includes perturbed background, translation jitter, rotation, and scaling.
+Table 10: Train/test statistics for ModelNet10, ModelNet40, ShapeNet and ScanObjectNN.
+ModelNet10 ModelNet40 ShapeNet ScanObjectNN
+# of training samples 3991 9843 35708 11416
+# of test samples 908 2468 10261 2882
+# of classes 10 40 55 15
+# of points for each sample 1024 1024 1024 1024
+Network Architecture.
+We evaluate our method on five representative point-based architectures: PointNet [pointnet], PointNet++ [pointnet2], DGCNN [dgcnn], PointConv [pointconv], and Point Transformer [pointtransformer].
+  • PointNet is a widely used neural network designed for processing 3D point clouds.
+We use PointNet as the backbone in dataset distillation model.
+While the original PointNet includes two transformation modules for aligning input and features, we only use the input transformation module.
+The rest of the architecture follows the standard PointNet design.
+  • PointNet++ is used for cross-architecture evaluation.
+To reduce computational cost and accelerate training, we decrease the width of all MLP layers by half, while keeping the overall structure unchanged.
+We use two set abstraction modules with multi-scale grouping and one with single-scale grouping to extract point features, followed by two linear layers and a final classifier for prediction.
+  • DGCNN employs dynamic graph construction and EdgeConv layers to capture local geometric relationships by recomputing a $k$-NN graph ($k=20$) at each layer in the feature space.
+We use the default architecture, which consists of four EdgeConv layers followed by two linear layers and a final classifier.
+  • PointConv extends convolution to point clouds by accounting for non-uniform point densities during feature learning.
+The architecture consists of three density-aware set abstraction layers, followed by two linear layers and a final classifier.
+  • Point Transformer applies self-attention mechanisms to point clouds to capture both local and global dependencies.
+We use the default configuration with four transformer blocks, followed by two linear layers and a final classifier.
+Implementation Details.
+We optimized the synthetic dataset $\boldsymbol{\mathcal{S}}$ using stochastic gradient descent (SGD) with a learning rate of 0.01, a momentum of 0.5, a weight decay of 0, and a batch size of 8 per class sampled from the original dataset $\boldsymbol{\mathcal{T}}$, while the batch size of the synthetic dataset was set equal to the number of synthetic samples per class (PPC).
+The synthetic dataset optimization was performed for 1,500 iterations, and the corresponding configuration is summarized in Table 11.
+To preserve fine-grained geometric details, the SADM loss was computed using the feature maps before the max pooling layer.
+Additionally, a secondary loss term was applied to the top-1 sorted feature values in each channel to emphasize semantically dominant regions.
+The loss weights $\lambda_{1}$ and $\lambda_{2}$ were determined based on PPC, and set to 0.002, 0.006, and 0.02 for $\lambda_{1}$, and 0.001, 0.003, and 0.01 for $\lambda_{2}$ when PPC was 1, 3, and 10, respectively.
+The rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ were jointly optimized with the synthetic dataset using SGD with a momentum of 0.5 and a weight decay of 0.
+The learning rates were set to 0.5 for $\theta_{x}$ and $\theta_{z}$, and 5.0 for $\theta_{y}$, reflecting that the samples in datasets are vertically aligned.
+A step decay scheduler was used with a step size of 100 and a decay factor of 0.5.
+This setup is detailed in Table 11.
+For evaluation, we trained all backbone networks including PointNet, PointNet++, DGCNN, PointConv, and Point Transformer using SGD with a learning rate of 0.01, a momentum of 0.9, and a weight decay of 0.0005.
+A step decay scheduler was applied with a step size of 250 and a decay factor of 0.1.
+Each model was trained for 500 epochs with a batch size of 8.
+These test-time settings are listed in Table 11.
+All experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
+Table 11: Hyperparameter settings used for (a) optimizing the synthetic dataset, (b) optimizing the rotation parameters, and (c) evaluation network.
+Hyperparameters PPC1 PPC3 PPC10
+Optimizer SGD SGD SGD
+Momentum 0.5 0.5 0.5
+Weight Decay 0.0 0.0 0.0
+Learning Rate 0.01 0.01 0.01
+Iteration 1500 1500 1500
+$\lambda_{1}$ 0.002 0.006 0.02
+$\lambda_{2}$ 0.001 0.003 0.01
+Batch Size $(\boldsymbol{\mathcal{T}})$ 8 8 8
+Hyperparameters
+Optimizer SGD
+Momentum 0.5
+Weight Decay 0.0
+Learning Rate $(\theta_{x})$ 0.5
+Learning Rate $(\theta_{y})$ 5.0
+Learning Rate $(\theta_{z})$ 0.5
+Scheduler StepLR
+Step Size 100
+Gamma 0.5
+Hyperparameters
+Optimizer SGD
+Momentum 0.9
+Weight Decay 0.0005
+Learning Rate 0.01
+Epochs 500
+Batch Size 8
+Scheduler StepLR
+Step Size 250
+Gamma 0.1
+Figure 4: Visualization of synthetic samples of ModelNet40 [modelnet40] at PPC 1, obtained by using the random selection [herding(icarl)](top), DataDAM [datadam](middle), and MTT [mtt1](bottom).
+Baselines.
+All baseline methods were implemented on 3D point clouds using PointNet as the feature extractor, following their original designs.
+All methods were trained for 1,500 iterations using a batch size of 8 per class from the original dataset $\boldsymbol{\mathcal{T}}$, and the training configurations are summarized in Table 12.
+Details for each method are as follows
+  • DC [gm]: We followed the original framework and performed gradient matching. The number of inner and outer loop steps were set to 1/1/10 and 1/5/3 for PPC values of 1, 3, and 10, respectively
+  • DM [dm]: Following the original setup, we used the feature vector obtained before the final classifier to compute the matching loss between the original and synthetic datasets.
+  • MTT [mtt1]: We followed the original framework for trajectory matching, with modifications only in the hyperparameter settings.
+  • DataDAM [datadam]: We followed the original framework for feature matching, making only minor adjustments to hyperparameters.
+Table 12: Hyperparameter settings of the baselines.
+DC DM MTT DataDAM
+Mode PointNet PointNet PointNet PointNet
+Learning Rate 0.0001 1 0.0001 0.0001
+Batch Size $(\boldsymbol{\mathcal{T}})$ 8 8 8 8
+Iteration 1500 1500 1500 1500
+Inner Loop 1 / 1 / 10
+Outer Loop 1 / 5 / 3
+Appendix C Comparison with Additional Baselines
+Table 13 presents the additional comparison results of the classification accuracy with DataDAM [datadam] and MTT [mtt1].
+For DataDAM and MTT, we used a possible learning rate to prevent divergence on the synthetic datasets.
+The performance of both DataDAM and MTT is almost similar to that of random selection [herding(icarl)], as shown in the second and third rows in Figure 4 where most points remain nearly unchanged from their random initialization in the first row.
+Table 13: Classification accuracy of the proposed method compared with DataDAM [datadam] and MTT [mtt1] initialized with randomly selected samples.
+Datasets ModelNet10 ModelNet40 ShapeNet ScanObjectNN
+PPC 1 3 10 1 3 10 1 3 10 1 3 10
+MTT 28.95 76.73 85.19 34.42 59.81 73.88 33.90 52.92 62.73 13.94 20.28 34.07
+DataDAM 40.51 76.60 86.56 34.54 60.37 74.94 35.70 54.56 63.64 16.29 20.51 35.45
+Ours 44.70 84.96 87.79 55.80 72.08 80.07 50.20 63.74 68.35 17.29 31.84 43.91
+Appendix D Additional Ablation Studies
+To further support the effectiveness of the proposed rotation optimization method, we compared widely used rotation augmentation that applies random rotations to the training data to address the rotation variation of 3D point clouds.
+While this strategy is effective when dealing with datasets showing severe rotational variations, such as ScanObjectNN, it does not achieve good performance on datasets where the 3D objects are fully or partially aligned, such as ModelNet10, ShapeNet, and ModelNet40.
+As shown in Table 14, the proposed learnable rotation estimation method consistently improves the performance across all datasets, however, the random rotation augmentation introduces unnecessary variation in already aligned datasets, degrading the performance.
+This highlights the benefit of explicitly modeling the orientation via optimization over the naive augmentation.
+Table 14: Performance comparison between the proposed rotation optimization and the random rotation augmentation (Aug.). All experiments were conducted at PPC 3.
+Prop. Aug. Aligned Mixed Rotated
+_ _ 65.01 59.96 20.42
+_ 58.92 52.56 31.42
+_ 74.35 72.08 31.84
+Appendix E Additional Qualitative Results
+Figure 5 shows randomly sampled 3D models from ModelNet40.
+Figures 6, 7, and 8 visualize the synthesized models at PPC 1 obtained by using DC, DM, and the proposed method, respectively.
+While DC and DM maintain the geometric shapes of the original data with only slight movement of some points, the proposed method generates noise-free samples with high visual quality across all classes.
+Figure 9 compares the optimization process for DM [dm] and the proposed method.
+DM fails to converge even after multiple training iterations, however the proposed method progressively refines the point cloud models into more structured shapes.
+Figure 10 shows the results of the proposed method on ModelNet40 at different PPC values: 1, 3, and 10.
+As PPC increases, the distilled synthetic datasets exhibit more diverse shapes and orientations.
+In particular, as shown in the last row, the human class appears in diverse poses.
+Additional qualitative results of the proposed method on the ShapeNet and ScanObjectNN datasets can be found in Figures 11 and 12, respectively.
+Figure 5: Randomly sampled 3D models of all classes from ModelNet40 [modelnet40].
+Figure 6: The synthetic dataset distilled from ModelNet40 by using DC at PPC 1.
+Figure 7: The synthetic dataset distilled from ModelNet40 by using DM at PPC 1.
+Figure 8: The synthetic dataset distilled from ModelNet40 by using the proposed method at PPC 1.
+Figure 9: Optimization process in the ModelNet40 at PPC 1 initialized from uniform noise, distilled by using (a) DM [dm] and (b) the proposed method.
+Figure 10: The synthetic dataset distilled from ModelNet40 by using the proposed method at three PPC values of 1, 3, and 10, respectively.
+Figure 11: The synthetic dataset distilled from ShapeNet by using the proposed method at PPC 1.
+Figure 12: The synthetic dataset distilled from ScanObjectNN by using the proposed method at PPC 1.